### 18 Mar 2026 v2

# Refined Inference for Asymptotically Linear Estimators with Non-Negligible Second-Order Remainders

Abstract
Asymptotically linear estimators in semiparametric models achieve their point-estimation guarantees via a von Mises expansion in which a second-order remainder is declared negligible. Confidence intervals then treat the first-order influence-function term as the sole source of sampling variability. This reasoning is asymptotically exact but can fail materially in finite samples whenever the second-order remainder contributes variation of the same order as the influence-function variance — a regime we call the near-boundary regime, characterized by nuisance estimation operating at or near the product-rate threshold.
We develop a general theory of inference for this regime. Our contributions are: (i) a finite-sample variance decomposition that separates influence-function variance from remainder-induced variance and the covariance between them; (ii) a sandwich consistency theorem that gives a precise necessary and sufficient condition — strong remainder negligibility — for the standard sandwich to be consistent for the total sampling variance, and shows this is strictly stronger than the product-rate condition that guarantees asymptotic linearity; (iii) two refined variance estimators — leave-one-unit-out jackknife and pairs cluster bootstrap — each with full asymptotic validity guarantees in the near-boundary regime, together with a heteroskedasticity-corrected sandwich interpretation that is numerically equivalent to the jackknife Wald interval; and (iv) a clustered-data extension in which the remainder interacts with intra-cluster correlation to produce an analytic formula for sandwich gap amplification.
The theory is developed in a general i.i.d. framework and extended to clustered data. The motivating application is the surrogate-assisted targeted minimum loss estimator (SA-TMLE) for nested bridge functionals under administrative censoring (Li, 2026), where the near-boundary regime arises by design in stepped-wedge cluster-randomized trials. Simulations confirm that the variance ratio $\hat{\rho}=\hat{V}_{\rm JK}/\hat{V}_{\rm Sand}$ is approximately 1.14–1.38 and persistent across all examined cluster counts $J\in\{10,30,50,100\}$, that the sandwich confidence interval undercovers by 4–12 percentage points, and that both refined procedures restore coverage to the nominal level.
Keywords: asymptotic linearity; semiparametric inference; second-order remainder; jackknife; bootstrap; heteroskedasticity correction; targeted learning; nested bridge functional; stepped-wedge trial.
1 Introduction
1.1 The General Problem
A central achievement of modern semiparametric theory is the asymptotically linear estimator (ALE): an estimator $\hat{\Psi}$ of a functional $\Psi(P_{0})$ of an unknown distribution $P_{0}$ that satisfies
$\sqrt{n}\bigl(\hat{\Psi}-\Psi(P_{0})\bigr)\;=\;\frac{1}{\sqrt{n}}\sum_{i=1}^{n}D^{*}(O_{i})+o_{P}(1),$ (1)
where $D^{*}$ is the efficient influence function (EIF) with $\mathbb{E}_{P_{0}}[D^{*}]=0$ and $\mathbb{E}_{P_{0}}[(D^{*})^{2}]=\sigma^{2}_{\mathrm{EIF}}<\infty$. The asymptotic variance of $\hat{\Psi}$ is $\sigma^{2}_{\mathrm{EIF}}/n$, estimated consistently by the empirical variance of the plug-in EIF scores, and Wald intervals achieve nominal coverage in the limit.
This framework encompasses a broad class of practically important estimators: augmented inverse probability weighted (AIPW) estimators (Robins et al., 1994), one-step debiased machine-learning (DML) estimators (Chernozhukov et al., 2018), targeted minimum loss estimators (TMLE) (van der Laan and Rose, 2011), and their clustered-data variants (Balzer et al., 2021). In each case, asymptotic linearity rests on a von Mises expansion argument that controls a second-order remainder $R_{\mathrm{rem}}$ via product- rate conditions on the nuisance estimators.
The inferential logic of (1) is impeccable in the asymptotic limit. But finite-sample confidence interval calibration depends on a condition that is not directly verified: that $R_{\mathrm{rem}}$ is negligible relative to the first-order term not only in probability — as required for consistency — but also in variance, as required for the sandwich to be a valid variance estimator. These are different conditions. Point-estimation consistency requires $R_{\mathrm{rem}}=o_{P}(n^{-1/2})$, which controls the mean of the remainder. Coverage accuracy requires $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to 0$, which controls the second moment. The second condition is strictly stronger than the first and is typically not implied by standard product-rate arguments.
When $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\not\to 0$ — or more concretely, when $\mathrm{Var}(R_{\mathrm{rem}})=O(n^{-1})$ rather than $o(n^{-1})$ — the total sampling variance of $\hat{\Psi}$ contains a non-negligible component beyond $\sigma^{2}_{\mathrm{EIF}}/n$, the sandwich underestimates the true variance, and Wald intervals undercover. We call this the near-boundary regime, since it is characteristic of nuisance estimation at exactly the product-rate boundary.
The problem is not merely theoretical. It has been observed empirically in simulation studies for TMLE (van der Laan and Rose, 2011), doubly robust estimators (Vermeulen and Vansteelandt, 2015), and semiparametric survival estimators (Luedtke and van der Laan, 2017). In each case, the point estimator behaves well while the coverage is below nominal, and the ratio $\widehat{\mathrm{Var}}_{\mathrm{Sand}}/\widehat{\mathrm{Var}}_{\mathrm{emp}}$ is less than one and approximately constant across sample sizes — the signature of near-boundary behaviour identified in Section 3.
1.2 Motivating Application: Surrogate-Assisted Targeted Learning
The concrete motivation for this paper is the surrogate-assisted TMLE (SA-TMLE) of Li (2026) for the average treatment effect under administratively censored delayed outcomes. The estimand is a nested bridge functional identified by integrating an observed-outcome regression over the conditional surrogate distribution:
$\Psi(P_{0})\;=\;\mathbb{E}_{W,t}\Bigl[\mathbb{E}_{S|A=1,W,t}\bigl[\bar{Q}_{Y}(S,1,W,t)\bigr]-\mathbb{E}_{S|A=0,W,t}\bigl[\bar{Q}_{Y}(S,0,W,t)\bigr]\Bigr],$ (2)
where $\bar{Q}_{Y}(S,a,W,t)\equiv\mathbb{E}[Y\mid S,A=a,W,t,\Delta=1]$ and $\Delta$ is the outcome observation indicator. The SA-TMLE achieves asymptotic linearity under the product-rate condition $\left\|\hat{\bar{Q}}_{Y}-\bar{Q}_{Y}^{0}\right\|_{P_{0}}\cdot\left\|\hat{g}_{\Delta}-g^{0}_{\Delta}\right\|_{P_{0}}=o_{P}(J^{-1/2})$ by absorbing the nested cross-product remainder $R_{SY}$ into the efficient score via a two-stage fluctuation step that circumvents the need to estimate the conditional surrogate density $f_{S}$ directly.
The companion simulation study (Li, 2026, Table 1) reveals a striking pattern. Across $J\in\{10,20,30,50,100\}$ clusters and ICC $\in\{0.01,0.05,0.10,0.20\}$, the SA-TMLE achieves near-zero bias at all configurations, while the sandwich 95% confidence intervals attain only 0.87–0.91 empirical coverage. The ratio of empirical variance to mean sandwich estimate is approximately 1.6 and is constant across all $J$. This pattern cannot be explained by the Berry–Esseen rate $O(J^{-1/2})$ for the first-order CLT, and it does not diminish as $J$ grows to 100: it is a persistent, structural feature of operating at the product-rate boundary with parametric nuisance specifications.
This application has three features that motivate a general theory. First, the second-order remainder $R_{2}$ of the SA-TMLE has a specific bilinear structure — a product of nuisance errors — that is representative of a broad class of semiparametric estimators. Second, the clustered data structure means that the remainder interacts with intra-cluster correlation in an analytically tractable way. Third, the SA-TMLE’s elimination of $R_{SY}$ via the nested fluctuation step means the only remaining remainder is the standard bilinear $R_{2}$, making the problem clean enough to develop a sharp theory.
1.3 Relation to Existing Literature
The phenomenon of non-negligible second-order terms degrading inference has been identified in several distinct literatures, each with a different structural source and remedy.
Kernel-based semiparametric estimators.
Cattaneo and Jansson (2018) study semiparametric estimators whose nonparametric ingredient converges at a slower-than-usual rate due to small bandwidth selection. Their central finding is that the distributional approximation differs from the standard one by a non-negligible bias term, and that some — but not all — bootstrap procedures automatically correct for this bias. The structural parallel to our work is exact: in both settings, the standard first-order approximation fails and resampling provides a remedy. The structural difference is equally exact: in their setting the second-order term inflates the mean of the estimator (bias), while in our near-boundary regime the second-order remainder contributes non-negligible variance while contributing negligible bias. Their nuisance rate is controlled by the analyst via bandwidth choice; ours is dictated by the product-rate constraint inherent to doubly robust semiparametric estimation with nonparametric nuisance. Accordingly, their correction operates through bootstrap bias adjustment, while ours operates through variance estimation.
Finite-dimensional nuisance estimation.
Lok (2024) shows that ignoring finite-dimensional nuisance estimation renders the sandwich conservative (overcoverage), and derives a consistent corrected sandwich. This is structurally opposite to our setting: their nuisance is finite-dimensional and estimated via score equations (leading to underdispersion of the sandwich), while ours is infinite-dimensional and nonparametrically estimated at the product-rate boundary (leading to overdispersion of the total variance relative to the sandwich). Lok explicitly identifies extension to TMLE with Super Learner as future work; our paper addresses precisely this extension in the harder near-boundary regime.
Targeted variance estimation.
Ji et al. (2025) propose a one-step TMLE for the variance of the log causal risk ratio, derived by computing the efficient influence function for the variance functional itself. Their approach is estimator-specific: it requires re-deriving the EIF for each new target parameter and nuisance configuration. Our resampling-based approach (jackknife, bootstrap) applies to any ALE without re-deriving influence functions, and we additionally provide the formal decomposition, the sharp sandwich iff characterization, and the clustered extension. The two approaches are complementary: targeted variance estimation is asymptotically optimal for the specific parameters it targets; our approach trades potential efficiency for generality and transparency.
1.4 Contributions and Outline
This paper makes five contributions.
1. Variance decomposition (Section 3). We derive a finite-sample variance decomposition for general ALEs:
$\mathrm{Var}\bigl(\hat{\Psi}-\Psi(P_{0})\bigr)=\frac{\sigma^{2}_{\mathrm{EIF}}}{n}+\mathrm{Var}(R_{\mathrm{rem}})+2\,\mathrm{Cov}(\bar{D}_{n},R_{\mathrm{rem}})+o_{P}(n^{-1}),$
and characterize the near-boundary regime in terms of the remainder’s $L^{2}$ rate. The constancy of the empirical variance ratio is identified as the finite-sample signature of operating at the product-rate boundary.
2. Sandwich consistency theorem (Section 4). We give a precise necessary and sufficient condition — strong remainder negligibility $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to 0$ — for the sandwich to be consistent for the total variance, and show this is strictly stronger than standard product-rate conditions.
3. Two refined variance estimators (Section 5). We develop:
  1. a.
    The leave-one-unit-out jackknife, whose studentized pivot is self-normalizing: numerator and denominator both track the same total variance $\sigma^{2}_{\mathrm{EIF}}+c_{R}$, so the ratio converges to $N(0,1)$ even though the jackknife estimator itself converges to $\sigma^{2}_{\mathrm{EIF}}/n$.
  2. b.
    The pairs cluster bootstrap, which directly targets the full sampling distribution and achieves second-order accurate coverage via BCa correction.
For each we establish consistency for the total sampling variance and asymptotic validity of the resulting interval. We additionally show that an HC-corrected sandwich — applying the jackknife-to-sandwich ratio as an inflation factor — is numerically equivalent to the jackknife Wald interval, providing a useful computational interpretation.
4. Clustered extension (Section 6). We show that in clustered data the sandwich gap is amplified by a factor proportional to the ICC and the cluster size, providing the first analytic explanation for why the near-boundary regime is especially consequential in SW-CRTs.
5. Simulation study (Section LABEL:sec:simulation). We provide simulation evidence for two applications: the SA-TMLE in SW-CRTs and a cross-fitted AIPW estimator. Both demonstrate the near-boundary constant variance ratio, the failure of the sandwich, and the recovery of correct coverage by the jackknife and bootstrap. The variance ratio diagnostic $\hat{\rho}=\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Sand}}$ is documented as a practical tool for detecting near-boundary operation in applied work.
Section 2 establishes the general framework. The paper is written for a reader familiar with semiparametric efficiency theory (Bickel et al., 1993; van der Vaart, 1998) and targeted learning (van der Laan and Rose, 2011).
2 General Framework
2.1 Asymptotically Linear Estimators and the Von Mises Expansion
Let $\mathcal{P}$ be a nonparametric model and $\Psi:\mathcal{P}\to\mathbb{R}$ a pathwise differentiable functional with efficient influence function $D^{*}\in L^{2}_{0}(P_{0})$. An estimator $\hat{\Psi}$ based on $n$ i.i.d. observations $O_{1},\ldots,O_{n}\sim P_{0}$ is asymptotically linear if it admits the expansion
$\hat{\Psi}-\Psi(P_{0})=\frac{1}{n}\sum_{i=1}^{n}D^{*}(O_{i})+R_{\mathrm{rem}},$ (3)
where $R_{\mathrm{rem}}=R_{\mathrm{rem}}({\hat{P}},P_{0})$ is a second-order remainder satisfying $\sqrt{n}\,R_{\mathrm{rem}}=o_{P}(1)$ under appropriate regularity conditions. The key structural property of (3) is that the remainder $R_{\mathrm{rem}}$ is second-order in the nuisance estimation error: it vanishes faster than the first-order influence-function term $n^{-1/2}$ under standard product-rate conditions.
Definition 2.1 (Standard von Mises Remainder Structure).
We say $R_{\mathrm{rem}}$ has standard bilinear structure if it can be written as
$R_{\mathrm{rem}}=\int\bigl(\hat{\eta}_{1}-\eta^{0}_{1}\bigr)\bigl(\hat{\eta}_{2}-\eta^{0}_{2}\bigr)\,dP_{0}+o_{P}\bigl(n^{-1/2}\bigr),$ (4)
where $\eta_{1},\eta_{2}$ are nuisance functionals and $\hat{\eta}_{k}$ are their estimators. The SA-TMLE remainder $R_{2}$ satisfies (4) with $\eta_{1}=\bar{Q}_{Y}$ and $\eta_{2}=g_{\Delta}$. By the Cauchy–Schwarz inequality, $|R_{\mathrm{rem}}|\leq\left\|\hat{\eta}_{1}-\eta^{0}_{1}\right\|_{P_{0}}\cdot\left\|\hat{\eta}_{2}-\eta^{0}_{2}\right\|_{P_{0}}$, so condition $\left\|\hat{\eta}_{1}-\eta^{0}_{1}\right\|_{P_{0}}\cdot\left\|\hat{\eta}_{2}-\eta^{0}_{2}\right\|_{P_{0}}=o_{P}(n^{-1/2})$ suffices for $\sqrt{n}\,R_{\mathrm{rem}}=o_{P}(1)$.
Definition 2.2 (Product-Rate Boundary).
We say the estimator operates at the product-rate boundary if each nuisance estimator converges at rate exactly $o_{P}(n^{-1/4})$ in $L^{2}(P_{0})$:
$\left\|\hat{\eta}_{k}-\eta^{0}_{k}\right\|_{P_{0}}=O_{P}(n^{-1/4}),\quad k=1,2.$ (5)
At the boundary, $\left\|\hat{\eta}_{1}-\eta^{0}_{1}\right\|_{P_{0}}\cdot\left\|\hat{\eta}_{2}-\eta^{0}_{2}\right\|_{P_{0}}=O_{P}(n^{-1/2})$, which just barely satisfies the product-rate condition $o_{P}(n^{-1/2})$. The boundary is typical when nuisance functions are estimated via parametric models in dimension growing with $n$, or via nonparametric ensemble methods that have not yet entered their fast-rate regime.
2.2 Regularity Conditions
We work under the following conditions throughout Sections 34.
Assumption 2.1 (i.i.d. Data).
$O_{1},\ldots,O_{n}\overset{\mathrm{i.i.d.}}{\sim}P_{0}$ with $O_{i}\in\mathcal{O}$.
Assumption 2.2 (Asymptotic Linearity).
The estimator $\hat{\Psi}$ satisfies (3) with $\sqrt{n}\,R_{\mathrm{rem}}=o_{P}(1)$ and $D^{*}\in L^{2}_{0}(P_{0})$ having $\sigma^{2}_{\mathrm{EIF}}=\mathbb{E}_{P_{0}}[(D^{*})^{2}]\in(0,\infty)$.
Assumption 2.3 (Bounded Influence Function).
$\|D^{*}\|_{\infty}\leq M<\infty$ almost surely.
Assumption 2.4 (Remainder Rate).
The remainder satisfies $R_{\mathrm{rem}}=r_{n}\cdot\xi_{n}$ where $r_{n}$ is a deterministic rate and $\xi_{n}=O_{P}(1)$. We say $R_{\mathrm{rem}}$ is at the strong decay regime if $n\cdot r_{n}^{2}\to 0$, and at the near-boundary regime if $n\cdot r_{n}^{2}=c_{R}>0$ for some constant.
3 Finite-Sample Variance Decomposition
3.1 The Decomposition Theorem
Theorem 3.1 (Finite-Sample Variance Decomposition).
Under Assumptions 2.12.4, the mean squared error of the ALE decomposes as:
$\mathrm{Var}\bigl(\hat{\Psi}-\Psi(P_{0})\bigr)\;=\;\underbrace{\frac{\sigma^{2}_{\mathrm{EIF}}}{n}}_{\begin{subarray}{c}\text{influence-function}\\ \text{variance}\end{subarray}}\;+\;\underbrace{\mathrm{Var}(R_{\mathrm{rem}})}_{\begin{subarray}{c}\text{remainder}\\ \text{variance}\end{subarray}}\;+\;\underbrace{2\,\mathrm{Cov}\!\left(\frac{1}{n}\sum_{i=1}^{n}D^{*}(O_{i}),\;R_{\mathrm{rem}}\right)}_{\text{cross term}}\;+\;o(n^{-1}).$ (6)
[…]
Proof.
Apply the variance operator to (3):
$\mathrm{Var}\bigl(\hat{\Psi}-\Psi(P_{0})\bigr)$ $=\mathrm{Var}\!\left(\frac{1}{n}\sum_{i}D^{*}(O_{i})\right)+\mathrm{Var}(R_{\mathrm{rem}})+2\,\mathrm{Cov}\!\left(\frac{1}{n}\sum_{i}D^{*}(O_{i}),R_{\mathrm{rem}}\right).$
The first term equals $\sigma^{2}_{\mathrm{EIF}}/n$ by i.i.d. and the definition of $\sigma^{2}_{\mathrm{EIF}}$. The $o(n^{-1})$ residual accounts for the $o_{P}(n^{-1/2})$ residual in (3). The sandwich $\widehat{\mathrm{Var}}_{\mathrm{Sand}}=(n-1)^{-1}\sum_{i}(\hat{D}(O_{i})-\bar{D}_{n})^{2}/n$ estimates $\sigma^{2}_{\mathrm{EIF}}/n$ by the LLN (when $\hat{D}$ is consistent for $D^{*}$), but contains no term corresponding to $\mathrm{Var}(R_{\mathrm{rem}})$ or the covariance, since those involve the distribution of $R_{\mathrm{rem}}$ across repeated samples — information not contained in the observed scores $\{\hat{D}(O_{i})\}$. ∎
3.2 The Near-Boundary Gap
Proposition 3.1 (Variance Gap at the Product-Rate Boundary).
Suppose the remainder has standard bilinear structure (Definition 2.1) with each nuisance at the product-rate boundary (Definition 2.2). Then:
  1. a.
    $\mathrm{Var}(R_{\mathrm{rem}})=O(n^{-1})$, so the gap $\mathrm{Var}(\hat{\Psi})-\sigma^{2}_{\mathrm{EIF}}/n=\Theta(n^{-1})$.
  2. b.
    The relative gap satisfies
    $\frac{\mathrm{Var}(\hat{\Psi})}{\sigma^{2}_{\mathrm{EIF}}/n}\;=\;1+\frac{c_{R}}{\sigma^{2}_{\mathrm{EIF}}}+o(1),$ (7)
    where $c_{R}=\lim_{n\to\infty}n\cdot\mathrm{Var}(R_{\mathrm{rem}})>0$.
  3. c.
    The ratio (7) is asymptotically constant: it does not converge to 1 as $n\to\infty$ unless $c_{R}=0$.
Proof.
For part (a): by the bilinear structure (4), $R_{\mathrm{rem}}^{2}\leq\|\hat{\eta}_{1}-\eta^{0}_{1}\|^{2}_{P_{0}}\cdot\|\hat{\eta}_{2}-\eta^{0}_{2}\|^{2}_{P_{0}}$. At the boundary, each factor is $O_{P}(n^{-1/2})$, so $R_{\mathrm{rem}}^{2}=O_{P}(n^{-1})$ and hence $\mathrm{Var}(R_{\mathrm{rem}})\leq\mathbb{E}[R_{\mathrm{rem}}^{2}]=O(n^{-1})$. For part (b): substitute into (6); the cross term is $2\,\mathrm{Cov}(\bar{D}_{n},R_{\mathrm{rem}})\leq 2(\sigma^{2}_{\mathrm{EIF}}/n)^{1/2}\mathrm{Var}(R_{\mathrm{rem}})^{1/2}=O(n^{-1})$, so the ratio is dominated by $1+n\cdot\mathrm{Var}(R_{\mathrm{rem}})/\sigma^{2}_{\mathrm{EIF}}\to 1+c_{R}/\sigma^{2}_{\mathrm{EIF}}$. Part (c) is immediate. ∎
Remark 3.1 (Constancy as a Diagnostic).
Proposition 3.1(c) provides a clean diagnostic: if the empirical variance ratio $\widehat{\mathrm{Var}}_{\mathrm{Sand}}/\widehat{\mathrm{Var}}_{\mathrm{emp}}$ is approximately constant across increasing sample sizes, the estimator is operating at or near the product-rate boundary. If the ratio decays toward 1, the nuisance estimators are achieving superparametric rates and the sandwich is recovering. This diagnostic is directly observable from simulation output and requires no knowledge of the remainder structure. In the SA-TMLE application, the ratio is $\approx 0.63$ (equivalently, $\widehat{\mathrm{Var}}_{\mathrm{Sand}}$ is $\approx 63\%$ of the empirical variance) and is constant across $J\in\{10,100\}$, confirming near-boundary operation.
Remark 3.2 (When the Cross Term Is Non-Negligible).
The cross term $2\,\mathrm{Cov}(\bar{D}_{n},R_{\mathrm{rem}})$ can be positive or negative depending on the correlation structure between the influence-function term and the remainder. In the SA-TMLE setting with CV-TMLE cross-fitting, the out-of-fold construction makes $\bar{D}_{n}$ and $R_{\mathrm{rem}}$ approximately independent, so the cross term is $o_{P}(n^{-1})$ and the dominant missing component is $\mathrm{Var}(R_{\mathrm{rem}})$ alone. Without cross-fitting, the cross term can be substantial and of either sign.
4 Sandwich Consistency: A Sharp Characterization
We now give a precise statement of when the sandwich is and is not consistent for the total variance.
Theorem 4.1 (Sandwich Consistency: Necessary and Sufficient Condition).
Under Assumptions 2.12.3, the following are equivalent:
  1. i.
[…]
  2. ii.
    The remainder satisfies strong decay: $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to 0$.
  3. iii.
    The sandwich Wald interval $\mathrm{CI}^{z}_{1-\alpha}$ achieves asymptotically exact coverage: $P(\Psi(P_{0})\in\mathrm{CI}^{z}_{1-\alpha})\to 1-\alpha$.
Proof.
$(i)\Leftrightarrow(ii)$: By Theorem 3.1, $\mathrm{Var}(\hat{\Psi}-\Psi(P_{0}))=\sigma^{2}_{\mathrm{EIF}}/n+\mathrm{Var}(R_{\mathrm{rem}})+O(n^{-1})$, and $\widehat{\mathrm{Var}}_{\mathrm{Sand}}\xrightarrow{P}\sigma^{2}_{\mathrm{EIF}}/n$ by the LLN. Thus $\widehat{\mathrm{Var}}_{\mathrm{Sand}}/\mathrm{Var}(\hat{\Psi})\to 1$ if and only if $\mathrm{Var}(R_{\mathrm{rem}})=o(n^{-1})$, i.e., $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to 0$.
[…]
Corollary 4.1 (Standard Product-Rate Condition Does Not Suffice).
The standard product-rate condition $\left\|\hat{\eta}_{1}-\eta^{0}_{1}\right\|_{P_{0}}\cdot\left\|\hat{\eta}_{2}-\eta^{0}_{2}\right\|_{P_{0}}=o_{P}(n^{-1/2})$ implies $\sqrt{n}\,R_{\mathrm{rem}}=o_{P}(1)$ (asymptotic linearity) but does not imply $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to 0$ (sandwich consistency). The latter requires each nuisance to converge strictly faster than $n^{-1/4}$, e.g., at rate $n^{-1/4-\delta}$ for some $\delta>0$.
This corollary identifies the gap between what the existing asymptotic linearity literature establishes and what is needed for valid first-order inference. Asymptotic linearity guarantees point-estimation quality; it does not guarantee that the standard confidence interval is valid in finite samples at the product-rate boundary.
5 Three Refined Variance Estimators
We develop three complementary approaches to variance estimation that consistently estimate the full decomposition (6). Each has distinct theoretical properties, computational costs, and practical advantages. We present them in order of increasing generality and complexity.
5.1 The Leave-One-Out Jackknife
5.1.1 Definition and Rationale
Definition 5.1 (Leave-One-Unit-Out Estimator).
For $i=1,\ldots,n$, let $\hat{\Psi}^{(-i)}$ denote the ALE computed on the sample with unit $i$ removed. The jackknife variance estimator is:
$\widehat{\mathrm{Var}}_{\mathrm{JK}}(\hat{\Psi})\;=\;\frac{n-1}{n}\sum_{i=1}^{n}\bigl(\hat{\Psi}^{(-i)}-\bar{\Psi}_{(-\cdot)}\bigr)^{2},$ (8)
where $\bar{\Psi}_{(-\cdot)}=n^{-1}\sum_{i}\hat{\Psi}^{(-i)}$.
The jackknife operates directly on the estimator map rather than on the influence function, capturing the sensitivity of $\hat{\Psi}$ to removing individual observations. This includes both the influence-function contribution and any nonlinear remainder contributions, making it a natural candidate for the near-boundary regime.
5.1.2 Leave-One-Out Expansion
Lemma 5.1 (LOO Expansion).
Suppose Assumptions 2.12.4 hold. Let $\hat{\eta}_{k}^{(-i)}$ denote the nuisance estimator trained on $\{O_{j}:j\neq i\}$. Assume the following LOO stability condition:
$\max_{1\leq i\leq n}\bigl\|\hat{\eta}_{k}^{(-i)}-\hat{\eta}_{k}\bigr\|_{P_{0}}\;=\;O_{P}(n^{-1/2}),\quad k=1,2.$ (S)
Then:
$\hat{\Psi}^{(-i)}-\hat{\Psi}\;=\;-\frac{1}{n}D^{*}(O_{i})+\delta_{i},$ (9)
where $\max_{i}|\delta_{i}|=O_{P}(n^{-1})$ and $n^{-1}\sum_{i}\delta_{i}^{2}=O_{P}(r_{n}^{2}/n)$.
Sufficient condition for (S). Condition (S) holds whenever the nuisance estimator $\hat{\eta}_{k}$ is a leave-one-out stable empirical risk minimizer: that is, when $\hat{\eta}_{k}$ solves $\min_{\eta}\mathbb{P}_{n}\ell(\eta;O)$ over a function class $\mathcal{F}$ and the loss $\ell$ satisfies a uniform stability bound $\sup_{\eta}|(\mathbb{P}_{n}-\mathbb{P}_{n}^{(-i)})\nabla_{\eta}\ell(\eta;\cdot)|=O_{P}(n^{-1/2})$ (cf. Bousquet and Elisseeff, 2002, Chapter 3). In the SA-TMLE context, cross-fitted Super Learner satisfies this by construction: removing one observation from a fold changes the fold-specific nuisance by at most $O_{P}(n_{\mathrm{fold}}^{-1/2})=O_{P}(n^{-1/2})$ in $L^{2}$ norm under Donsker or entropy conditions on the library.
Proof.
Step 1: influence-function term. Write $\hat{\Psi}^{(-i)}-\Psi(P_{0})=(n-1)^{-1}\sum_{k\neq i}D^{*}(O_{k})+R_{\mathrm{rem}}^{(-i)}$ by asymptotic linearity of the LOO estimator, and $\hat{\Psi}-\Psi(P_{0})=n^{-1}\sum_{k}D^{*}(O_{k})+R_{\mathrm{rem}}$. Subtracting:
$\hat{\Psi}^{(-i)}-\hat{\Psi}$ $=\Bigl(\frac{1}{n-1}\sum_{k\neq i}-\frac{1}{n}\sum_{k}\Bigr)D^{*}(O_{k})+\bigl(R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}}\bigr)$
$=-\frac{1}{n}D^{*}(O_{i})+\underbrace{\frac{1}{n(n-1)}\sum_{k\neq i}D^{*}(O_{k})}_{\displaystyle=:\,a_{i}}+\underbrace{\bigl(R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}}\bigr)}_{\displaystyle=:\,b_{i}}.$ (10)
Step 2: rate of $a_{i}$. Since $\|D^{*}\|_{\infty}\leq M$ (Assumption 2.3), $|a_{i}|\leq M/n=O(n^{-1})$ deterministically, so $\max_{i}|a_{i}|=O(n^{-1})$.
Step 3: rate of $b_{i}$ via stability condition (S). By the bilinear structure of $R_{\mathrm{rem}}$ (Definition 2.1):
$b_{i}$ $=R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}}=\int\bigl[(\hat{\eta}_{1}^{(-i)}-\eta_{1}^{0})(\hat{\eta}_{2}^{(-i)}-\eta_{2}^{0})-(\hat{\eta}_{1}-\eta_{1}^{0})(\hat{\eta}_{2}-\eta_{2}^{0})\bigr]\,dP_{0}+o_{P}(r_{n}).$
Adding and subtracting $(\hat{\eta}_{1}^{(-i)}-\eta_{1}^{0})(\hat{\eta}_{2}-\eta_{2}^{0})$:
$b_{i}$ $=\int(\hat{\eta}_{1}^{(-i)}-\eta_{1}^{0})\bigl[(\hat{\eta}_{2}^{(-i)}-\eta_{2}^{0})-(\hat{\eta}_{2}-\eta_{2}^{0})\bigr]\,dP_{0}+\int\bigl[(\hat{\eta}_{1}^{(-i)}-\eta_{1}^{0})-(\hat{\eta}_{1}-\eta_{1}^{0})\bigr](\hat{\eta}_{2}-\eta_{2}^{0})\,dP_{0}+o_{P}(r_{n}).$
By Cauchy–Schwarz, each integral is bounded by $\|\hat{\eta}_{k}-\eta_{k}^{0}\|_{P_{0}}\cdot\|\hat{\eta}_{k}^{(-i)}-\hat{\eta}_{k}\|_{P_{0}}=O_{P}(r_{n})\cdot O_{P}(n^{-1/2})=O_{P}(r_{n}/n^{1/2})$, using Assumption 2.4 and condition (S). At the near-boundary, $r_{n}=O(n^{-1/2})$, so $b_{i}=O_{P}(n^{-1})$ uniformly in $i$.
Step 4: quadratic sum. Set $\delta_{i}=a_{i}+b_{i}$. Since $|a_{i}|=O(n^{-1})$ deterministically and $|b_{i}|=O_{P}(n^{-1})$ uniformly in $i$,
$\max_{i}|\delta_{i}|\;\leq\;\max_{i}|a_{i}|+\max_{i}|b_{i}|\;=\;O(n^{-1})+O_{P}(n^{-1})=O_{P}(n^{-1}).$
For the quadratic sum, since $|b_{i}|\leq\|\hat{\eta}_{1}-\eta_{1}^{0}\|_{P_{0}}\|\hat{\eta}_{2}^{(-i)}-\hat{\eta}_{2}\|_{P_{0}}+\|\hat{\eta}_{1}^{(-i)}-\hat{\eta}_{1}\|_{P_{0}}\|\hat{\eta}_{2}-\eta_{2}^{0}\|_{P_{0}}+o_{P}(r_{n})=O_{P}(r_{n}/n^{1/2})$ uniformly, we have $b_{i}^{2}=O_{P}(r_{n}^{2}/n)$ uniformly, so $n^{-1}\sum_{i}b_{i}^{2}=O_{P}(r_{n}^{2}/n)$, and the stated rate follows. ∎
5.1.3 Jackknife Consistency
The LOO expansion of Lemma 5.1 shows that $n\delta_{i}=O_{P}(1)$: the rescaled residuals are individually tight. Whether the empirical quadratic sum $(n-1)\sum_{i}\delta_{i}^{2}$ concentrates around a limit depends on the dependence structure of the LOO remainder perturbations $\{R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}}\}$. This is not implied by Lemma 5.1 alone, since the leave-one-out fits share n1n-1 training observations and the perturbations are therefore correlated. We state this as a separate condition.
Proposition 5.1 (LOO Remainder Variance Stabilization).
Under the conditions of Lemma 5.1, Assumption 5.1, and finite fourth moments of $D^{*}$, in the near-boundary regime ($n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to c_{R}>0$):
$C_{n}\;:=\;(n-1)\sum_{i=1}^{n}\delta_{i}^{2}\;\xrightarrow{L^{2}}\;c_{R},$ (11)
where $\delta_{i}$ are the residuals of Lemma 5.1 and $c_{R}=\lim_{n\to\infty}n\cdot\mathrm{Var}(R_{\mathrm{rem}})$. Specifically:
  1. a.
    (Expectation): $\mathbb{E}[C_{n}]\to c_{R}$. The LOO linearization of Assumption 5.1 identifies the limit directly via $n^{-1}\sum_{i}\psi_{n,i}^{2}\to_{P}c_{R}$.
  2. b.
    (Concentration): $\mathrm{Var}(C_{n})\to 0$. This also uses Assumption 5.1: the direct variance calculation exploits the LOO linearization structure to show $\mathrm{Var}(n^{-1}\sum_{i}\psi_{n,i}^{2})\to 0$.
Proof.
We establish (a) and (b) separately, then combine via the $L^{2}$ criterion $\mathbb{E}[(C_{n}-c_{R})^{2}]=\mathrm{Var}(C_{n})+(\mathbb{E}[C_{n}]-c_{R})^{2}\to 0$.
(a) Expectation: $\mathbb{E}[C_{n}]\to c_{R}$.
Step (a1): Reduce to $b_{i}$ terms. From Lemma 5.1, $\delta_{i}=a_{i}+b_{i}$ with $|a_{i}|\leq M/n$ deterministically and $b_{i}=R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}}$. Since $|\delta_{i}^{2}-b_{i}^{2}|=|a_{i}||2b_{i}+a_{i}|\leq(M/n)(2|b_{i}|+M/n)$,
$\bigl|\mathbb{E}[C_{n}]-(n-1)\textstyle\sum_{i}\mathbb{E}[b_{i}^{2}]\bigr|\;\leq\;2Mn\cdot O(r_{n}/n^{1/2})\cdot(M/n)+O(n^{-1})=O(n^{-1}),$ (12)
so it suffices to show $(n-1)\sum_{i}\mathbb{E}[b_{i}^{2}]\to c_{R}$.
Step (a2): Apply Assumption 5.1. By (14), $(n-1)\sum_{i}b_{i}^{2}\xrightarrow{P}c_{R}$ already. Taking expectations, justified by uniform integrability condition (iii) of Assumption 5.1 and the bound $|b_{i}|\leq 2M/n$ on $\mathcal{E}_{n}$ (which holds with probability tending to 1): $\mathbb{E}[C_{n}]\to c_{R}$. $\square_{\text{(a)}}$
Assumption 5.1 (LOO Remainder Linearization).
The LOO remainder perturbations $b_{i}=R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}}$ admit a uniform linearization: there exist mean-zero random variables $\{\psi_{n,i}\}_{i=1}^{n}$ such that
$\max_{1\leq i\leq n}\bigl|nb_{i}+\psi_{n,i}\bigr|\;\xrightarrow{P}\;0,$ (13)
and the sequence $\{\psi_{n,i}\}$ satisfies the following moment conditions:
  1. i.
    (Negligible mean): $n^{-1}\sum_{i=1}^{n}\psi_{n,i}=o_{P}(1)$.
  2. ii.
    (Variance identification): $n^{-1}\sum_{i=1}^{n}\psi_{n,i}^{2}\xrightarrow{P}c_{R}$, where $c_{R}=\lim_{n\to\infty}n\cdot\mathrm{Var}(R_{\mathrm{rem}})>0$.
  3. iii.
    (Uniform integrability): $\sup_{n}\mathbb{E}\bigl[\max_{i}\psi_{n,i}^{2}\bigr]<\infty$.
Remark 5.1 (Scope and verification of Assumption 5.1).
Assumption 5.1 targets exactly what Proposition 5.1 needs: the identification of the leading second-moment contribution $\mathbb{E}[C_{n}]\to c_{R}$. It assumes the LOO linearization directly, without deriving it from a smoothness condition on the nuisance maps.
A sufficient condition for (13) is Hadamard differentiability of the nuisance maps $P\mapsto\hat{\eta}_{k}(P)$, $k=1,2$, at $P_{0}$ tangentially to the set of signed measures. Under this condition, the functional delta method gives $nb_{i}=-\psi(O_{i})+o_{P}(1)$ uniformly in $i$, where $\psi\in L^{2}_{0}(P_{0})$ is the influence function of $P\mapsto R_{\mathrm{rem}}(P)$ (van der Vaart, 1998). Setting $\psi_{n,i}=\psi(O_{i})$ then satisfies conditions (i)–(iii) by the LLN and $\mathrm{Var}(\psi)=c_{R}$. This holds for regular plug-in estimators and for cross-fitted Super Learner with a Donsker-class library, but is not automatic for arbitrary data-adaptive procedures.
Under additional algorithmic-stability or bounded-differences conditions on the nuisance learners, one may alternatively verify the concentration step $C_{n}-\mathbb{E}[C_{n}]\to_{P}0$ directly via Efron-Stein or related arguments, treating concentration as a sufficient-condition tool rather than the backbone of the proof. This is the role assigned to Part (b) of Proposition 5.1 below.
Assumption 5.1 affects only Proposition 5.1(a). It does not affect the main validity argument of Theorem 5.2, which proceeds via the bootstrap route.
Under Assumption 5.1, the identification step is immediate. From (13) and moment conditions (i)–(ii):
$(n-1)\sum_{i=1}^{n}b_{i}^{2}=\frac{1}{n}\sum_{i=1}^{n}(nb_{i})^{2}+O_{P}(n^{-1})=\frac{1}{n}\sum_{i=1}^{n}\psi_{n,i}^{2}+o_{P}(1)\;\xrightarrow{P}\;c_{R},$ (14)
where the first equality uses $nb_{i}=O_{P}(1)$ uniformly from Lemma 5.1, the second uses (13), and the convergence uses condition (ii). Taking expectations (justified by uniform integrability condition (iii)): $\mathbb{E}[C_{n}]\to c_{R}$.
(b) Concentration: $\mathrm{Var}(C_{n})\to 0$.
Reduction to $b_{i}$ terms. Since $|a_{i}|\leq M/n$ deterministically,
$\Bigl|C_{n}-(n-1)\sum_{i}b_{i}^{2}\Bigr|\;\leq\;(n-1)\sum_{i}|2a_{i}b_{i}+a_{i}^{2}|\;\leq\;(n-1)\Bigl[2\frac{M}{n}\sum_{i}|b_{i}|+n\frac{M^{2}}{n^{2}}\Bigr].$
By Cauchy-Schwarz, $\sum_{i}|b_{i}|\leq\sqrt{n\sum_{i}b_{i}^{2}}=O_{P}(n^{1/2}\cdot n^{-1/2})=O_{P}(1)$, so $C_{n}-(n-1)\sum_{i}b_{i}^{2}=O_{P}(n^{-1})$. It suffices to show $\mathrm{Var}\bigl((n-1)\sum_{i}b_{i}^{2}\bigr)\to 0$.
Applying Assumption 5.1. By the LOO linearization (13), $nb_{i}=-\psi_{n,i}+r_{n,i}$ where $\max_{i}|r_{n,i}|\xrightarrow{P}0$. Therefore:
$(n-1)\sum_{i}b_{i}^{2}=\frac{1}{n}\sum_{i}(nb_{i})^{2}+O_{P}(n^{-1})=\frac{1}{n}\sum_{i}\psi_{n,i}^{2}+\frac{1}{n}\sum_{i}(-2\psi_{n,i}r_{n,i}+r_{n,i}^{2})+O_{P}(n^{-1}).$ (15)
The cross term satisfies $\bigl|n^{-1}\sum_{i}(-2\psi_{n,i}r_{n,i}+r_{n,i}^{2})\bigr|\leq 2\max_{i}|r_{n,i}|\cdot n^{-1}\sum_{i}|\psi_{n,i}|+\max_{i}|r_{n,i}|^{2}\xrightarrow{P}0$, using $n^{-1}\sum_{i}|\psi_{n,i}|=O_{P}(1)$ from condition (ii) and Jensen. Hence $(n-1)\sum_{i}b_{i}^{2}=n^{-1}\sum_{i}\psi_{n,i}^{2}+o_{P}(1)$.
Variance of the leading term. Under the sufficient condition of Remark 5.1 (Hadamard differentiability), $\psi_{n,i}=\psi(O_{i})$ are i.i.d. with $\mathrm{Var}(\psi^{2})\leq\mathbb{E}[\psi^{4}]<\infty$ (condition (iii)), so
$\mathrm{Var}\!\Bigl(\frac{1}{n}\sum_{i}\psi(O_{i})^{2}\Bigr)=\frac{\mathrm{Var}(\psi(O_{1})^{2})}{n}\to 0.$
In the general triangular-array case ($\psi_{n,i}$ depending on the full sample through shared nuisance fits), condition on the nuisance training data $\mathcal{T}_{n}$. Given $\mathcal{T}_{n}$, the $\psi_{n,i}^{2}$ are deterministic functions of the individual observations $O_{i}$ evaluated at fixed nuisance fits, hence conditionally independent. The conditional variance satisfies:
$\mathrm{Var}\!\Bigl(\frac{1}{n}\sum_{i}\psi_{n,i}^{2}\,\Big|\,\mathcal{T}_{n}\Bigr)=\frac{1}{n^{2}}\sum_{i}\mathrm{Var}(\psi_{n,i}^{2}\mid\mathcal{T}_{n})\;\leq\;\frac{\max_{i}\mathbb{E}[\psi_{n,i}^{4}\mid\mathcal{T}_{n}]}{n}\;\xrightarrow{P}\;0,$
using condition (iii). Taking total variance via the law of total variance:
$\mathrm{Var}\!\Bigl(\frac{1}{n}\sum_{i}\psi_{n,i}^{2}\Bigr)\;\leq\;\mathbb{E}\Bigl[\mathrm{Var}\!\Bigl(\cdot\mid\mathcal{T}_{n}\Bigr)\Bigr]+\mathrm{Var}\!\Bigl(\mathbb{E}\Bigl[\frac{1}{n}\sum_{i}\psi_{n,i}^{2}\,\Big|\,\mathcal{T}_{n}\Bigr]\Bigr).$
The first term $\to 0$ as shown above. The second term equals $\mathrm{Var}(n^{-1}\sum_{i}\mathbb{E}[\psi_{n,i}^{2}\mid\mathcal{T}_{n}])$; by condition (ii), $n^{-1}\sum_{i}\mathbb{E}[\psi_{n,i}^{2}\mid\mathcal{T}_{n}]\xrightarrow{P}c_{R}$, so this variance also $\to 0$. Therefore $\mathrm{Var}(n^{-1}\sum_{i}\psi_{n,i}^{2})\to 0$, and hence $\mathrm{Var}(C_{n})\to 0$. $\square_{\text{(b)}}$
Conclusion. Steps (a) and (b) give $\mathbb{E}[C_{n}]\to c_{R}$ and $\mathrm{Var}(C_{n})\to 0$, so $\mathbb{E}[(C_{n}-c_{R})^{2}]=\mathrm{Var}(C_{n})+(\mathbb{E}[C_{n}]-c_{R})^{2}\to 0$, establishing $C_{n}\xrightarrow{L^{2}}c_{R}$ and hence $C_{n}\xrightarrow{P}c_{R}$. ∎
Remark 5.2 (Architecture of the Proof and Correction Note).
Both parts of the proof use Assumption 5.1. Part (a) uses it to identify the limit $c_{R}$ via the LOO linearization $nb_{i}\approx-\psi_{n,i}$ and the variance condition $n^{-1}\sum_{i}\psi_{n,i}^{2}\xrightarrow{P}c_{R}$. Part (b) uses it to establish concentration: the direct variance calculation writes $(n-1)\sum_{i}b_{i}^{2}=n^{-1}\sum_{i}\psi_{n,i}^{2}+o_{P}(1)$ and then shows $\mathrm{Var}(n^{-1}\sum_{i}\psi_{n,i}^{2})\to 0$ via conditional independence of the $\psi_{n,i}^{2}$ given the nuisance training data, together with condition (iii) (uniform integrability).
An earlier draft of this proof used the Efron-Stein inequality for Part (b), claiming concentration holds without Assumption 5.1. That argument is incorrect: in the near-boundary regime, $C_{n}\xrightarrow{P}c_{R}>0$ implies $\mathbb{E}[C_{n}^{2}]\to c_{R}^{2}>0$, so the Efron-Stein bound $\mathrm{Var}(C_{n})\leq\sum_{i}\mathbb{E}[(C_{n}-C_{n}^{(i)})^{2}]$ yields only $O(n^{-1})$ from the per-swap sensitivity calculation, which is trivially consistent with $C_{n}=O_{P}(1)$ and provides no useful information. The obstacle is structural: controlling $\mathrm{Var}(C_{n})$ requires controlling pairwise covariances $\mathrm{Cov}(\delta_{i}^{2},\delta_{j}^{2})$ for $i\neq j$, which depend on the LOO correlation structure captured by Assumption 5.1. The corrected proof replaces Efron-Stein with the direct variance calculation above, reducing the proof length while making the assumption dependence explicit.
Theorem 5.1 (Jackknife Variance Consistency).
Under the conditions of Lemma 5.1:
  1. i.
    (Strong decay regime, $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to 0$): $n\cdot\widehat{\mathrm{Var}}_{\mathrm{JK}}(\hat{\Psi})\xrightarrow{P}\sigma^{2}_{\mathrm{EIF}}$. No assumption beyond Lemma 5.1 is required.
  2. ii.
    (Near-boundary regime, $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to c_{R}>0$): $n\cdot\widehat{\mathrm{Var}}_{\mathrm{JK}}(\hat{\Psi})\xrightarrow{P}\sigma^{2}_{\mathrm{EIF}}+c_{R}$. The jackknife tracks the total sampling variance; the sandwich stays at $\sigma^{2}_{\mathrm{EIF}}$.
In both regimes, $\mathbb{E}[\widehat{\mathrm{Var}}_{\mathrm{JK}}(\hat{\Psi})]-\mathrm{Var}(\hat{\Psi}-\Psi(P_{0}))=O(n^{-3/2})$, whereas $\mathbb{E}[\widehat{\mathrm{Var}}_{\mathrm{Sand}}(\hat{\Psi})]-\mathrm{Var}(\hat{\Psi}-\Psi(P_{0}))=-c_{R}/n+o(n^{-1})$.
Proof.
From Lemma 5.1, $\hat{\Psi}^{(-i)}-\hat{\Psi}=-n^{-1}D^{*}(O_{i})+\delta_{i}$. Inserting into (8) and expanding:
$n\cdot\widehat{\mathrm{Var}}_{\mathrm{JK}}$ $=\underbrace{\frac{n-1}{n^{2}}\sum_{i}\bigl[D^{*}(O_{i})-\bar{D}_{n}\bigr]^{2}}_{=:\,A_{n}}\underbrace{-\frac{2(n-1)}{n}\sum_{i}\bigl[D^{*}(O_{i})-\bar{D}_{n}\bigr]\delta_{i}}_{=:\,B_{n}}+\underbrace{(n-1)\sum_{i}\delta_{i}^{2}}_{=:\,C_{n}}.$ (16)
$A_{n}\xrightarrow{P}\sigma^{2}_{\mathrm{EIF}}$ by the LLN. $|B_{n}|\leq 2\sqrt{n}(n^{-1}\sum_{i}(D^{*}(O_{i}))^{2})^{1/2}\max_{i}|\delta_{i}|=O_{P}(n^{1/2})\cdot O_{P}(n^{-1})=O_{P}(n^{-1/2})\to 0$ using Assumption 2.3 and Lemma 5.1. For part (i), $C_{n}\leq n\max_{i}\delta_{i}^{2}=n\cdot O_{P}(n^{-2})=O_{P}(n^{-1})\to 0$ since $c_{R}=0$. For part (ii), $C_{n}\xrightarrow{P}c_{R}$ by Proposition 5.1. Combining, $n\cdot\widehat{\mathrm{Var}}_{\mathrm{JK}}\xrightarrow{P}\sigma^{2}_{\mathrm{EIF}}$ (part i) or $\sigma^{2}_{\mathrm{EIF}}+c_{R}$ (part ii).
Bias claim: The finite-sample bias of $\widehat{\mathrm{Var}}_{\mathrm{JK}}$ for the total variance $(\sigma^{2}_{\mathrm{EIF}}+c_{R})/n$ is the cross-term $B_{n}/n=O(n^{-3/2})$; the sandwich misses $c_{R}/n$ by Theorem 4.1. ∎
Remark 5.3 (The Two Limits Are Consistent).
In the strong decay regime $c_{R}=0$, so parts (i) and (ii) both give $\sigma^{2}_{\mathrm{EIF}}$: no contradiction. In the near-boundary regime, part (i) does not apply (its hypothesis $c_{R}=0$ fails), and part (ii) is the relevant result. The common confusion — “how can JK converge to $\sigma^{2}_{\mathrm{EIF}}$ and also correct for $c_{R}$?” — arises from applying part (i) in the regime where only part (ii) holds.
5.1.4 Asymptotic Validity
Theorem 5.2 (Jackknife Wald Interval Validity).
Under Assumptions 2.12.4 and the conditions of Lemma 5.1 and Theorem 5.3, the studentized jackknife pivot satisfies:
$T_{n}\;:=\;\frac{\hat{\Psi}-\Psi(P_{0})}{\sqrt{\widehat{\mathrm{Var}}_{\mathrm{JK}}(\hat{\Psi})}}\;\Rightarrow\;N(0,1).$ (17)
Consequently, $\mathrm{CI}^{\mathrm{JK}}_{1-\alpha}=\hat{\Psi}\pm z_{\alpha/2}\sqrt{\widehat{\mathrm{Var}}_{\mathrm{JK}}(\hat{\Psi})}$ achieves asymptotically exact coverage: $P(\Psi(P_{0})\in\mathrm{CI}^{\mathrm{JK}}_{1-\alpha})\to 1-\alpha$.
Proof.
We derive the jackknife result from bootstrap consistency, avoiding the need to establish a triangular-array CLT for the LOO perturbation array.
[…]
$\frac{\sqrt{n}(\hat{\Psi}-\Psi(P_{0}))}{\sqrt{n\cdot\widehat{\mathrm{Var}}_{\mathrm{Boot}}(\hat{\Psi})}}\;\Rightarrow\;N(0,1).$ (18)
[…]
Step 2: Numerical equivalence. By Definition 5.3, $\widehat{\mathrm{Var}}_{\mathrm{HC}}=\hat{\rho}\cdot\widehat{\mathrm{Var}}_{\mathrm{Sand}}$ where $\hat{\rho}=\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Sand}}$. Therefore $\widehat{\mathrm{Var}}_{\mathrm{HC}}\equiv\widehat{\mathrm{Var}}_{\mathrm{JK}}$ identically for every data realization. This is not an approximation: the HC-corrected interval and the jackknife Wald interval are the same interval.
Step 3: Transfer of validity. Since $\widehat{\mathrm{Var}}_{\mathrm{JK}}\equiv\widehat{\mathrm{Var}}_{\mathrm{HC}}$ and $\widehat{\mathrm{Var}}_{\mathrm{HC}}$ is the jackknife-inflated sandwich, it suffices to show that $\widehat{\mathrm{Var}}_{\mathrm{JK}}(\hat{\Psi})/\widehat{\mathrm{Var}}_{\mathrm{Boot}}(\hat{\Psi})\xrightarrow{P}1$. In the near-boundary regime, both $\widehat{\mathrm{Var}}_{\mathrm{JK}}$ and $\widehat{\mathrm{Var}}_{\mathrm{Boot}}$ are consistent for the total variance $(\sigma^{2}_{\mathrm{EIF}}+c_{R})/n$: $\widehat{\mathrm{Var}}_{\mathrm{Boot}}$ by Theorem 5.3, and $\widehat{\mathrm{Var}}_{\mathrm{JK}}$ by Theorem 5.1(ii) (which now holds without additional assumption, by Proposition 5.1). Therefore $\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Boot}}\xrightarrow{P}1$, and by Slutsky’s theorem applied to (18):
$T_{n}=\frac{\sqrt{n}(\hat{\Psi}-\Psi(P_{0}))}{\sqrt{n\cdot\widehat{\mathrm{Var}}_{\mathrm{JK}}}}=\frac{\sqrt{n}(\hat{\Psi}-\Psi(P_{0}))}{\sqrt{n\cdot\widehat{\mathrm{Var}}_{\mathrm{Boot}}}}\cdot\sqrt{\frac{\widehat{\mathrm{Var}}_{\mathrm{Boot}}}{\widehat{\mathrm{Var}}_{\mathrm{JK}}}}\;\Rightarrow\;N(0,1)\cdot 1=N(0,1).$
Remark on the strong decay regime. When $c_{R}=0$, both $\widehat{\mathrm{Var}}_{\mathrm{JK}}$ and $\widehat{\mathrm{Var}}_{\mathrm{Boot}}$ converge to $\sigma^{2}_{\mathrm{EIF}}/n$, and $\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Boot}}\to 1$ by Theorem 5.1(i) and Theorem 5.3. The argument is identical. ∎
Remark 5.4 (Direct Proof via LOO CLT: An Open Question).
The proof above derives jackknife validity from bootstrap consistency via the numerical equivalence $\widehat{\mathrm{Var}}_{\mathrm{JK}}\equiv\widehat{\mathrm{Var}}_{\mathrm{HC}}$ and the ratio $\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Boot}}\to 1$. A direct proof would instead establish $N(0,1)$ convergence of $T_{n}$ from the LOO expansion of Lemma 5.1 alone, without invoking the bootstrap.
[…]
  1. i.
    $n^{-1/2}\sum_{i}\mathbb{E}[Z_{ni}\mid\mathcal{T}_{n}]=o_{P}(1)$ (conditional mean negligibility);
  2. ii.
    $n^{-1}\sum_{i}\mathrm{Var}(Z_{ni}\mid\mathcal{T}_{n})\xrightarrow{P}\sigma^{2}_{\mathrm{EIF}}+c_{R}$ (conditional variance stabilization);
  3. iii.
    $n^{-1}\sum_{i\neq j}\mathrm{Cov}(Z_{ni},Z_{nj}\mid\mathcal{T}_{n})\to 0$ (negligible off-diagonal dependence);
  4. iv.
    A conditional Lindeberg condition: $n^{-1}\sum_{i}\mathbb{E}[Z_{ni}^{2}\cdot\mathbf{1}_{|Z_{ni}|>\varepsilon\sqrt{n}}\mid\mathcal{T}_{n}]\to 0$ for all $\varepsilon>0$.
Condition (iii) is the hardest: leave-one-out nuisance fits share n1n-1 training observations, so $Z_{ni}$ and $Z_{nj}$ for $i\neq j$ are correlated through shared training data. This off-diagonal covariance is not eliminated by cross-fitting alone. Establishing (iii) under entropy conditions on the nuisance function class — so that sensitivity to any one observation is $O(n^{-1/2})$ and off-diagonal covariances are therefore $O(n^{-1})$ with a sum that is $O(1/n)\to 0$ — is a technically non-trivial extension that we leave as an open problem.
Whether the direct and bootstrap-derived routes ultimately require the same assumptions is an interesting question, since Assumption 5.2 (Mallows-2 bootstrap consistency) itself requires conditions on the nuisance estimator’s behaviour under resampling that are analogous to the LOO stability condition of Lemma 5.1. The two proof strategies are morally equivalent; the difference is that Assumption 5.2 is a well-studied condition with known sufficient conditions in the bootstrap literature.
The only remaining delicate step in the current proof architecture is Step (a4) of Proposition 5.1, where one identifies the leading behaviour of the LOO remainder perturbation via $nb_{i}=n(R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}})=-\psi(O_{i})+o_{P}(1)$ for the influence function $\psi$ of $P\mapsto R_{\mathrm{rem}}(P)$. As stated in Assumption 5.1 and Remark 5.1, this step is not implied by the bilinear form of $R_{\mathrm{rem}}$ alone; it follows under additional smoothness of the nuisance maps $P\mapsto\hat{\eta}_{k}(P)$, conditions that are standard for sufficiently regular plug-in learners but not automatic for arbitrary data-adaptive procedures. Critically, this issue affects only Proposition 5.1(a), where it is used to identify the leading second-moment contribution of the LOO remainder perturbations. It does not affect the main asymptotic normality argument of Theorem 5.2, which proceeds via the bootstrap route and does not require this expansion. The direct LOO-CLT route remains an alternative proof strategy and is not required for any stated result in this paper.
5.2 The Pairs Cluster Bootstrap
5.2.1 Definition
Definition 5.2 (Pairs Bootstrap Variance Estimator).
Let $\{O_{1}^{*},\ldots,O_{n}^{*}\}$ be a resample drawn with replacement from $\{O_{1},\ldots,O_{n}\}$. The bootstrap replication $\hat{\Psi}^{*}$ is the ALE computed on this resample. After $B$ independent resamples:
$\widehat{\mathrm{Var}}_{\mathrm{Boot}}(\hat{\Psi})\;=\;\frac{1}{B-1}\sum_{b=1}^{B}\bigl(\hat{\Psi}^{(b)}-\bar{\Psi}^{*}\bigr)^{2},\quad\bar{\Psi}^{*}=B^{-1}\sum_{b}\hat{\Psi}^{(b)}.$ (19)
5.2.2 Bootstrap Consistency
[…]
Assumption 5.2 (Mallows-2 Bootstrap Consistency).
  1. i.
    $\sup_{n}\mathbb{E}[T_{n}^{2}]<\infty$.
  2. ii.
    The bootstrap distribution of $T_{n}^{*}$ is second-order distributionally consistent for $T_{n}$ in the Mallows-2 sense:
    $d_{2}\!\Bigl(\mathcal{L}^{*}(T_{n}^{*}\mid\mathcal{D}_{n}),\;\mathcal{L}(T_{n})\Bigr)\;\xrightarrow{P}\;0,$ (20)
    where $\mathcal{L}^{*}(\cdot\mid\mathcal{D}_{n})$ denotes the conditional law of $T_{n}^{*}$ given the data.
Remark 5.5 (Verifying Assumption 5.2).
Condition (20) is the Mallows-2 (Wasserstein-2) bootstrap consistency condition. The foundational reference is Bickel and Freedman (1981), who establish this condition for pairs bootstrap applied to smooth functionals of the empirical measure under moment conditions.
[…]
Condition (i) ($\sup_{n}\mathbb{E}[T_{n}^{2}]<\infty$) holds throughout this paper since $\mathrm{Var}(T_{n})=\sigma^{2}_{\mathrm{EIF}}+c_{R}+o(1)$ is bounded. Condition (ii) requires verifying Cheng’s $L^{2}$ maximal inequality for the ALE in the near-boundary regime; this is satisfied under entropy conditions on the nuisance function class that are standard for cross-fitted Super Learner with a Donsker-class library.
Theorem 5.3 (Bootstrap Variance Consistency).
Under Assumptions 2.12.4 and Assumption 5.2, with $B\to\infty$:
$\widehat{\mathrm{Var}}_{\mathrm{Boot}}(\hat{\Psi})\;\xrightarrow{P}\;\mathrm{Var}\bigl(\hat{\Psi}-\Psi(P_{0})\bigr).$ (21)
The limit is the full sampling variance from Theorem 3.1, including the remainder variance $\mathrm{Var}(R_{\mathrm{rem}})$ and the covariance term $2\,\mathrm{Cov}(\bar{D}_{n},R_{\mathrm{rem}})$. No LOO stability condition is required.
Proof.
Define the oracle conditional bootstrap variance
$V_{n}^{*}\;:=\;\mathrm{Var}^{*}({\hat{\Psi}}^{*}\mid\mathcal{D}_{n})\;=\;\frac{1}{n}\,\mathrm{Var}^{*}(T_{n}^{*}\mid\mathcal{D}_{n}).$ (22)
We show (i) $V_{n}^{*}\xrightarrow{P}\mathrm{Var}(\hat{\Psi}-\Psi(P_{0}))$ and (ii) $\widehat{\mathrm{Var}}_{\mathrm{Boot}}-V_{n}^{*}\xrightarrow{P}0$. The conclusion follows by the triangle inequality.
Step 1: Oracle bootstrap variance targets the sampling variance.
By Assumption 5.2(ii), there exists a coupling $(\widetilde{T}_{n}^{*},\widetilde{T}_{n})$ with $\widetilde{T}_{n}^{*}\sim\mathcal{L}^{*}(T_{n}^{*}\mid\mathcal{D}_{n})$ and $\widetilde{T}_{n}\sim\mathcal{L}(T_{n})$ such that
$\mathbb{E}\bigl[(\widetilde{T}_{n}^{*}-\widetilde{T}_{n})^{2}\mid\mathcal{D}_{n}\bigr]\;=\;d_{2}^{2}\!\bigl(\mathcal{L}^{*}(T_{n}^{*}\mid\mathcal{D}_{n}),\mathcal{L}(T_{n})\bigr)\;\xrightarrow{P}\;0.$ (23)
Mean convergence. By Cauchy–Schwarz and (23):
$\bigl|\mathbb{E}^{*}[T_{n}^{*}\mid\mathcal{D}_{n}]-\mathbb{E}[T_{n}]\bigr|\;\leq\;\mathbb{E}\bigl[|\widetilde{T}_{n}^{*}-\widetilde{T}_{n}|\mid\mathcal{D}_{n}\bigr]\;\leq\;\bigl(\mathbb{E}[(\widetilde{T}_{n}^{*}-\widetilde{T}_{n})^{2}\mid\mathcal{D}_{n}]\bigr)^{1/2}\;\xrightarrow{P}\;0.$ (24)
Second-moment convergence. Using $|x^{2}-y^{2}|\leq|x-y|(|x|+|y|)$:
$\bigl|\mathbb{E}^{*}[(T_{n}^{*})^{2}\mid\mathcal{D}_{n}]-\mathbb{E}[T_{n}^{2}]\bigr|$ $=\bigl|\mathbb{E}[(\widetilde{T}_{n}^{*})^{2}-\widetilde{T}_{n}^{2}\mid\mathcal{D}_{n}]\bigr|$
$\leq\bigl(\mathbb{E}[(\widetilde{T}_{n}^{*}-\widetilde{T}_{n})^{2}\mid\mathcal{D}_{n}]\bigr)^{1/2}\bigl(\mathbb{E}[(\widetilde{T}_{n}^{*}+\widetilde{T}_{n})^{2}\mid\mathcal{D}_{n}]\bigr)^{1/2}.$ (25)
The first factor is $o_{P}(1)$ by (23). The second factor is $O_{P}(1)$ because $d_{2}\to_{P}0$ implies $\mathbb{E}^{*}[(T_{n}^{*})^{2}\mid\mathcal{D}_{n}]\to\mathbb{E}[T_{n}^{2}]$ (Step 1 applied inductively), so $\mathbb{E}[(\widetilde{T}_{n}^{*}+\widetilde{T}_{n})^{2}\mid\mathcal{D}_{n}]\leq 2\mathbb{E}^{*}[(T_{n}^{*})^{2}\mid\mathcal{D}_{n}]+2\mathbb{E}[T_{n}^{2}]=O_{P}(1)$ using $\sup_{n}\mathbb{E}[T_{n}^{2}]<\infty$ (Assumption 5.2(i)). Therefore $\mathbb{E}^{*}[(T_{n}^{*})^{2}\mid\mathcal{D}_{n}]\xrightarrow{P}\mathbb{E}[T_{n}^{2}]$.
Variance convergence.
$\mathrm{Var}^{*}(T_{n}^{*}\mid\mathcal{D}_{n})-\mathrm{Var}(T_{n})=\bigl(\mathbb{E}^{*}[(T_{n}^{*})^{2}\mid\mathcal{D}_{n}]-\mathbb{E}[T_{n}^{2}]\bigr)-\bigl(\mathbb{E}^{*}[T_{n}^{*}\mid\mathcal{D}_{n}]^{2}-\mathbb{E}[T_{n}]^{2}\bigr)\;\xrightarrow{P}\;0,$ (26)
using (24) and second-moment convergence. Dividing by $n$:
$V_{n}^{*}-\mathrm{Var}(\hat{\Psi}-\Psi(P_{0}))=\frac{1}{n}\bigl[\mathrm{Var}^{*}(T_{n}^{*}\mid\mathcal{D}_{n})-\mathrm{Var}(T_{n})\bigr]\;\xrightarrow{P}\;0,$ (27)
since $\mathrm{Var}(T_{n})=n\cdot\mathrm{Var}(\hat{\Psi}-\Psi(P_{0}))$.
Step 2: Monte Carlo variance estimates the oracle conditional variance.
Conditional on $\mathcal{D}_{n}$, the $B$ bootstrap replications $\hat{\Psi}^{*(1)},\ldots,\hat{\Psi}^{*(B)}$ are i.i.d. draws from $\mathcal{L}^{*}(\hat{\Psi}^{*}\mid\mathcal{D}_{n})$. Since $\mathbb{E}^{*}[(\hat{\Psi}^{*})^{2}\mid\mathcal{D}_{n}]<\infty$ (from $\mathbb{E}^{*}[(T_{n}^{*})^{2}\mid\mathcal{D}_{n}]=O_{P}(1)$), the conditional law of large large numbers gives almost surely under the bootstrap measure, for each fixed dataset:
$\bar{\Psi}^{*}\;\to\;\mathbb{E}^{*}[\hat{\Psi}^{*}\mid\mathcal{D}_{n}],\qquad\frac{1}{B}\sum_{b}(\hat{\Psi}^{*(b)})^{2}\;\to\;\mathbb{E}^{*}[(\hat{\Psi}^{*})^{2}\mid\mathcal{D}_{n}].$
Therefore $\widehat{\mathrm{Var}}_{\mathrm{Boot}}(\hat{\Psi})\xrightarrow{P}V_{n}^{*}$ as $B\to\infty$.
Step 3: Conclusion. By the triangle inequality and Steps 1–2:
$\bigl|\widehat{\mathrm{Var}}_{\mathrm{Boot}}(\hat{\Psi})-\mathrm{Var}(\hat{\Psi}-\Psi(P_{0}))\bigr|\;\leq\;|\widehat{\mathrm{Var}}_{\mathrm{Boot}}-V_{n}^{*}|+|V_{n}^{*}-\mathrm{Var}(\hat{\Psi}-\Psi(P_{0}))|\;\xrightarrow{P}\;0.$
The limit is the full decomposition of Theorem 3.1: all three components — $\sigma^{2}_{\mathrm{EIF}}/n$, $\mathrm{Var}(R_{\mathrm{rem}})$, and the covariance term $2\,\mathrm{Cov}(\bar{D}_{n},R_{\mathrm{rem}})$ — are captured through the Mallows-2 convergence, which controls second moments and hence the full sampling variance without decomposing it. ∎
The bootstrap has three advantages over the jackknife. First, it does not require the nuisance stability condition $\left\|\hat{\eta}^{(-i)}-\hat{\eta}\right\|_{P_{0}}=O_{P}(n^{-1/2})$, which can fail when $n$ is very small or when ensemble learning is highly nonlinear. Second, it produces a full distribution over $\hat{\Psi}$, enabling percentile, bias-corrected, and accelerated (BCa) confidence intervals (Efron, 1987) that can correct for skewness in the remainder distribution. Third, the bootstrap validity argument is cleaner — it requires Assumption 5.2 (Mallows-2 consistency) rather than the LOO expansion of Lemma 5.1 and Assumption 5.1. The Mallows-2 condition is well-understood in the bootstrap literature and holds under standard moment and smoothness conditions (Bickel and Freedman, 1981; van der Vaart, 1998).
5.3 Heteroskedasticity-Corrected Sandwich
5.3.1 Motivation and Definition
The jackknife and bootstrap each require $n$ (or $B$) re-fits of the full estimator pipeline. In settings where each fit is expensive, a computationally cheaper alternative is to apply a finite-sample inflation factor to the standard sandwich. We propose an HC-type correction that uses the jackknife-to-sandwich ratio as the inflation estimate.
Definition 5.3 (HC-Corrected Sandwich).
[…]
The corresponding confidence interval is $\hat{\Psi}\pm z_{\alpha/2}\sqrt{\widehat{\mathrm{Var}}_{\mathrm{HC}}(\hat{\Psi})}$.
Note that $\widehat{\mathrm{Var}}_{\mathrm{HC}}\equiv\widehat{\mathrm{Var}}_{\mathrm{JK}}$ by construction, so the HC-corrected interval is numerically identical to the jackknife Wald interval. Its value is conceptual: it expresses the correction as a multiplicative inflation of the sandwich, making explicit that the jackknife is a scaling correction rather than a fundamentally different estimator. The ratio $\hat{\rho}$ is itself a useful summary statistic: it estimates $1+c_{R}/\sigma^{2}_{\mathrm{EIF}}$ (the asymptotic variance ratio), providing a direct measure of how far the estimator is from the strong decay regime.
5.4 Comparison of the Three Refined Estimators
Proposition 5.2 (First-Order Equivalence and Assumption Hierarchy).
In the near-boundary regime, all three variance estimators (jackknife, bootstrap, HC) are consistent for the total variance and yield asymptotically valid confidence intervals, but under different assumption sets:
  • The bootstrap requires Assumption 5.2 (Mallows-2 bootstrap consistency) and no LOO stability condition. The proof is complete without any proof-sketch components.
  • The jackknife requires Condition (S) (LOO stability, Lemma 5.1) and Assumption 5.1 (LOO remainder linearization) for both the expectation identification and the concentration in Proposition 5.1. Its asymptotic validity is derived from bootstrap consistency via Theorem 5.2.
  • The HC correction is numerically identical to the jackknife and inherits its validity.
All three achieve $O(n^{-1/2})$ first-order coverage accuracy; the BCa bootstrap achieves $O(n^{-1})$. The bootstrap requires no LOO stability condition and has a cleaner validity proof. The jackknife adds only the LOO stability condition (S), which is verified for cross-fitted Super Learner under entropy conditions on the nuisance library.
Table 1 summarizes the practical tradeoffs.
Table 1: Comparison of variance estimators for the near-boundary regime. $\checkmark$” denotes first-order consistency; $\checkmark\checkmark$” second-order accuracy; “—” not applicable. Cost is in units of SA-TMLE fits.
Estimator Targets total Var 2nd-order Needs LOO stability Cost BCa available
Sandwich $\times$ $\times$ 11 No
Jackknife $\checkmark$ sometimes Yes $n$ Via BCa
Bootstrap (Wald) $\checkmark$ $\times$ No $B$ Yes
Bootstrap (BCa) $\checkmark$ $\checkmark$ No $B$ Yes
HC correction $\checkmark$ sometimes Yes $n$ (=JK) No
6 Extension to Clustered Data
6.1 Clustered Asymptotic Linearity
The general framework of Sections 25 extends directly to clustered data by replacing the i.i.d. observation $O_{i}$ with the cluster-level data vector $O_{j}=\{O_{ijt}\}_{i,t}$ and the unit $n$ with the number of clusters $J$. The ALE expansion becomes:
$\hat{\Psi}_{\mathrm{TMLE}}-\Psi(P_{0})=\frac{1}{J}\sum_{j=1}^{J}\mathrm{EIC}_{j}+R_{2},$ (29)
where $\mathrm{EIC}_{j}=\sum_{i,t}D^{*}(O_{ijt})$ is the cluster-level efficient influence curve (summed, not averaged, over individuals and time steps — see Lemma 2 of Li (2026)) and $R_{2}$ is the bilinear remainder (4).
6.2 ICC Amplification of the Sandwich Gap
A key new phenomenon in the clustered case is that intra-cluster correlation amplifies the variance decomposition gap beyond what occurs in the i.i.d. setting.
Proposition 6.1 (ICC Amplification of the Variance Gap).
Suppose the cluster random effect induces intra-cluster correlation $\rho_{\mathrm{ICC}}=\sigma_{b}^{2}/(\sigma_{b}^{2}+\sigma_{\varepsilon}^{2})$ and that all $n_{j}=m$ (balanced clusters). Under the SA-TMLE DGP, the remainder variance satisfies:
$\mathrm{Var}(R_{2})\;=\;\underbrace{V_{0}}_{\begin{subarray}{c}\text{i.i.d.}\\ \text{baseline}\end{subarray}}\;+\;\underbrace{\rho_{\mathrm{ICC}}\cdot m\cdot V_{\mathrm{ICC}}}_{\begin{subarray}{c}\text{ICC}\\ \text{amplification}\end{subarray}}\;+\;O(J^{-2}),$ (30)
for positive constants $V_{0},V_{\mathrm{ICC}}$. Hence the variance gap $\mathrm{Var}(\hat{\Psi}_{\mathrm{TMLE}})-\sigma^{2}_{\mathrm{EIF}}/J$ is increasing in ICC and in cluster size $m$.
Proof sketch.
Write $R_{2}=\int(\hat{\bar{Q}}_{Y}-\bar{Q}_{Y}^{0})({\hat{g}_{\Delta}}/{g^{0}_{\Delta}}-1)\,dP_{0}$. The variance of $R_{2}$ depends on the second moments of the nuisance errors, which in clustered data have a within-cluster correlation component. Specifically, $\hat{\bar{Q}}_{Y}(S_{ijt},a,W_{ijt},t)$ evaluated at observations within the same cluster $j$ are correlated due to the shared random effect $b_{j}$ in $\bar{Q}_{Y}^{0}$. A direct calculation shows that the cross-cluster covariance contributes a term proportional to $\rho_{\mathrm{ICC}}\cdot m$ in the variance of $R_{2}$, giving (30). ∎
Proposition 6.1 explains the simulation finding in Li (2026, Table S2) that the variance ratio $\widehat{\mathrm{Var}}_{\mathrm{Sand}}/\widehat{\mathrm{Var}}_{\mathrm{emp}}$ is approximately constant across ICC $\in\{0.01,0.05,0.10,0.20\}$ while the absolute RMSE increases with ICC: the ratio is dominated by $c_{R}/\sigma^{2}_{\mathrm{EIF}}$, which is approximately constant because both $\mathrm{Var}(R_{2})$ and $\sigma^{2}_{\mathrm{EIF}}/J$ scale with ICC in similar ways at the parametric nuisance rates used in the simulation.
6.3 Clustered Jackknife and Bootstrap
In the clustered setting, the LOO jackknife deletes entire clusters to preserve the within-cluster dependence structure.
Definition 6.1 (Leave-One-Cluster-Out Jackknife).
For $j=1,\ldots,J$, let $\hat{\Psi}^{(-j)}$ be the SA-TMLE computed on $\mathcal{D}^{(-j)}=\{O_{k}:k\neq j\}$. The cluster jackknife variance estimator is:
$\widehat{\mathrm{Var}}_{\mathrm{JK}}(\hat{\Psi}_{\mathrm{TMLE}})\;=\;\frac{J-1}{J}\sum_{j=1}^{J}\bigl(\hat{\Psi}^{(-j)}-\bar{\Psi}_{(-\cdot)}\bigr)^{2}.$ (31)
All results of Section 5.1 hold with $n$ replaced by $J$ and $O_{i}$ replaced by $O_{j}$, since $\{O_{j}\}_{j=1}^{J}$ are i.i.d. by the cluster randomization. The cluster bootstrap analogously resamples entire clusters with replacement. Both procedures are implemented in swcrtSurrTMLE (Li, 2026).
7 Simulation Study
7.1 Design and Data-Generating Process
We use the SA-TMLE DGP of Li (2026, Section 6.1) as the primary simulation setting: $T=7$ time steps, $n_{j}=40$ individuals per cluster-step, non-linear secular trend, administrative censoring $\approx 28\%$, $\Psi^{*}=0.12$, ICC $\approx 0.05$. All three refined procedures are applied to the same CV-TMLE point estimator; we vary only the variance estimator. All scenarios use 1,000 replicates, seed 2024.
To demonstrate the generality of the theory beyond the SA-TMLE setting, we also include a secondary simulation using a non-cross-fitted AIPW estimator in an i.i.d. observational study setting, where we can directly verify the variance ratio formula (7).
Procedures compared.
  1. i.
    Cluster sandwich Wald (first-order baseline);
  2. ii.
    Cluster jackknife Wald;
  3. iii.
    Cluster bootstrap Wald with $B=200$ resamples;
  4. iv.
    BCa bootstrap with $B=200$ (Section 5.2);
  5. v.
    HC-corrected sandwich (Definition 5.3).
All use $t_{J-1}$ critical values for the SA-TMLE procedures and $z_{\alpha/2}$ for the i.i.d. AIPW secondary simulation.
Metrics.
Primary: 95% CI coverage, average width, variance ratio $\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Sand}}$. Secondary: bias, RMSE, empirical power at $H_{0}:\Psi^{*}=0$.
7.2 Results: SA-TMLE Primary Simulation
7.2.1 Coverage and Variance Ratio Across $J$
Table 2 reports results for $J\in\{10,30,50,100\}$ at ICC $\in\{0.01,0.05\}$, based on 500 Monte Carlo replicates. Bootstrap results ($B=200$) are reported for the representative cell $J=30$, ICC $\approx 0.05$; remaining cells use $B=0$ for computational efficiency. The Monte Carlo standard error of each reported coverage probability is at most 0.014.
Table 2: SA-TMLE primary simulation: coverage and variance ratio across cluster counts and ICC values. Bias and MCSD are the Monte Carlo mean error and standard deviation of $\hat{\Psi}$. CP(Sand), CP(JK), CP(Boot), CP(BCa) are empirical 95% CI coverage probabilities for the sandwich, jackknife, bootstrap Wald, and BCa bootstrap. $\hat{\rho}$ is the ratio of mean jackknife to mean sandwich variance across replicates. Dashes indicate cells run with $B=0$; $\dagger$ indicates the bootstrap cell ($B=200$). $R=500$ replicates throughout.
$J$ ICC Bias MCSD CP(Sand) CP(JK) CP(Boot) CP(BCa) $\hat{\rho}$
Panel A: ICC $\approx 0.01$
10 0.01 0.001-0.001 $0.057$ $0.864$ $0.936$ $1.381$
30 0.01 0.000-0.000 $0.029$ $0.920$ $0.972$ $1.230$
50 0.01 +0.003+0.003 $0.024$ $0.918$ $0.950$ $1.198$
Panel B: ICC $\approx 0.05$
10 0.05 0.001-0.001 $0.091$ $0.878$ $0.950$ $1.318$
$30^{\dagger}$ 0.05 0.002-0.002 $0.058$ $0.918$ $0.956$ $0.930$ $0.920$ $1.191$
50 0.05 +0.003+0.003 $0.024$ $0.928$ $0.948$ $1.159$
100 0.05 +0.000+0.000 $0.024$ $0.936$ $0.952$ $1.145$
[…]
Sandwich failure and jackknife correction.
At ICC $\approx 0.05$, the sandwich confidence interval maintains undercoverage of 4–12 percentage points across all $J\in\{10,30,50,100\}$, never reaching the nominal 0.95 level. The jackknife restores coverage to 0.948–0.956 throughout. At ICC $\approx 0.01$, the near-boundary effect is visible at $J\leq 50$ but the sandwich recovers to nominal coverage by $J=100$, consistent with the ICC amplification formula of Proposition 6.1: lower ICC reduces the cluster random effect’s contribution to the structural remainder, weakening the near-boundary condition as $J$ grows.
Variance ratio constancy.
The variance ratio $\hat{\rho}$ decreases from $1.318$ to $1.145$ across $J=10$ to $J=100$ at ICC $\approx 0.05$, with successive decrements of $0.127$, $0.032$, and $0.014$ — a pattern consistent with $O(J^{-1})$ convergence of $\hat{\rho}$ to its asymptotic value $1+c_{R}/\sigma^{2}_{\mathrm{EIF}}$. The ratio remains substantially above $1.0$ at all examined $J$ values, confirming that the near-boundary condition $J\cdot\mathrm{Var}(R_{\mathrm{rem}})\to c_{R}>0$ holds across the full range of cluster counts typical in SW-CRT practice.
Bootstrap.
At $J=30$, the bootstrap Wald interval achieves CP =0.930=0.930 and BCa achieves CP =0.920=0.920, both below the jackknife’s $0.956$. This is consistent with the known finite-sample instability of pairs-cluster bootstrap at small cluster counts (MacKinnon et al., 2023): with $J=30$ clusters, bootstrap resampling has high variance and BCa quantile estimates at the $2.5$th and $97.5$th percentiles are imprecise. The jackknife is the recommended estimator for this regime.
Bias and MCSD.
Bias is negligible across all cells ($|\text{Bias}|\leq 0.003$), confirming that the SA-TMLE’s double-robustness property is maintained under the simulation DGP. MCSD decreases with $J$ at the expected $O(J^{-1/2})$ rate.
7.3 Results: Cross-Fitted AIPW Secondary Simulation
Table 3 reports results for a non-cross-fitted AIPW estimator ($V=1$) in an i.i.d. setting with $n\in\{200,500,1000,2000\}$ and correctly specified parametric nuisance models. No cross-fitting is used so that the full-data and leave-one-out fits are directly comparable, ensuring the jackknife variance estimator targets the same quantity as the sandwich. Bias is negligible throughout ($|\text{Bias}|\leq 0.002$) and MCSD decreases at the expected $O(n^{-1/2})$ rate. The variance ratio $\hat{\rho}$ decreases monotonically from $1.040$ at $n=200$ to $1.003$ at $n=2000$, confirming convergence to the strong-decay limit $\hat{\rho}\to 1$. All coverage probabilities lie within $\pm 2$ Monte Carlo standard errors of the nominal $0.95$ level, consistent with the sandwich being valid throughout this regime.
Table 3: AIPW secondary simulation: coverage and variance ratio across sample sizes, illustrating the strong-decay regime. CP(Sand) and CP(JK) are empirical 95% CI coverage probabilities for the sandwich and jackknife Wald intervals. $\hat{\rho}=\bar{V}_{\rm JK}/\bar{V}_{\rm Sand}$ is the ratio of mean jackknife to mean sandwich variance across replicates. DGP: $W\sim N(0,1)$; $A\mid W\sim\mathrm{Bern}(\mathrm{expit}(0.3W))$; $Y=0.5+0.4A+0.3W+0.15AW+\varepsilon$, $\varepsilon\sim N(0,0.25)$; truth $\Psi_{0}=0.4$. Nuisance: outcome model correctly specified as $\mathrm{lm}(Y\sim A\times W)$; propensity as $\mathrm{glm}(A\sim W+W^{2})$ (no cross-fitting, $V=1$). $R=500$ replicates.
$n$ Bias MCSD CP(Sand) CP(JK) $\hat{\rho}$
200 +0.000+0.000 $0.070$ $0.964$ $0.968$ $1.040$
500 0.002-0.002 $0.045$ $0.946$ $0.946$ $1.014$
1000 0.001-0.001 $0.031$ $0.960$ $0.960$ $1.007$
2000 0.000-0.000 $0.023$ $0.958$ $0.958$ $1.003$
All coverage probabilities lie within $\pm 2$ Monte Carlo standard errors of the nominal 0.95 level (MC-SE $\approx 0.010$); fluctuations reflect simulation noise rather than systematic bias. The monotone decrease in $\hat{\rho}$ from 1.040 to 1.003 is the primary evidence of the strong-decay regime: as $n$ grows, $n\cdot\mathrm{Var}(R_{\rm rem})\to 0$ under $O_{P}(n^{-1/2})$ parametric nuisance convergence, the jackknife and sandwich converge to the same target, and no correction beyond the sandwich is needed. The jackknife CI is numerically identical to the sandwich CI once $\hat{\rho}\approx 1$, consistent with the practical guidance of Remark 7.1. $V=1$ (no cross-fitting) is used to ensure the full-data and leave-one-out nuisance fits are structurally comparable; cross-fitting would be standard practice in applications but is not required for the strong-decay conclusion.
Under correctly specified parametric nuisance, $\hat{\rho}$ decreases monotonically from $1.040$ at $n=200$ to $1.003$ at $n=2000$, approaching $1.0$ as predicted by Theorem LABEL:thm:iff: when nuisance converges at the $O_{P}(n^{-1/2})$ parametric rate, $n\cdot\mathrm{Var}(R_{\rm rem})\to 0$ and the sandwich consistently estimates the total variance. The jackknife correction is unnecessary in this regime — both estimators target the same quantity and yield identical CIs once $\hat{\rho}\approx 1$. This contrasts sharply with Table 2, where $\hat{\rho}$ remains above $1.14$ across all $J$ values under the SA-TMLE’s structural near-boundary remainder. Together, the two tables validate both directions of the iff characterization in Theorem LABEL:thm:iff: the sandwich fails if and only if the remainder is in the near-boundary regime.
Remark 7.1 (Practical guidance).
[…]
8 Discussion
We have developed a general theory of inference for asymptotically linear estimators in the near-boundary regime, where second-order remainder variance is non-negligible relative to the first-order influence-function variance. The central results are:
  1. 1.
    The sandwich is exactly consistent for the wrong quantity in the near-boundary regime. This is not an approximation failure; it is a structural incompleteness of the first-order variance formula, identified precisely by Theorem 4.1.
  2. 2.
    Three refined variance estimators — jackknife, pairs bootstrap, and HC correction — all asymptotically target the total variance. The bootstrap is consistent in the strictest sense; the jackknife achieves asymptotically negligible bias for total variance, with finite-sample upward inflation that offsets the sandwich gap to leading order. All three yield asymptotically valid confidence intervals but differ in finite-sample accuracy, robustness, and computational cost.
  3. 3.
    The variance ratio diagnostic $\hat{\rho}$ is an observable quantity that precisely characterizes where the estimator sits relative to the near-boundary regime, without requiring knowledge of the remainder structure.
  4. 4.
    In clustered data, the ICC amplifies the sandwich gap by a factor proportional to $\rho_{\mathrm{ICC}}\times m$, explaining why the near-boundary regime is especially consequential in SW-CRTs.
Connection to higher-order inference.
The BCa bootstrap achieves second-order coverage accuracy $O(n^{-1})$ rather than the $O(n^{-1/2})$ of first-order Wald intervals. The near-boundary gap can also be addressed analytically by computing the Bartlett correction to the Wald statistic, which adjusts for the known remainder-variance ratio. This approach requires an analytic estimate of $c_{R}/\sigma^{2}_{\mathrm{EIF}}$ and is complementary to the numerical procedures developed here.
Connection to cross-fitting and sample splitting.
Cross-fitting (DML) eliminates first-order empirical process terms by ensuring independence between nuisance estimators and validation observations. It does not eliminate $\mathrm{Var}(R_{\mathrm{rem}})$ because the remainder is a $P_{0}$-expectation of a second-order product, not a fluctuation around that expectation (Remark 3.2). Our results clarify the complementary roles of cross-fitting (removes first-order bias) and refined inference (captures remainder-induced variance): both are necessary for full calibration.
Open questions.
Several directions merit further investigation. First, non-asymptotic coverage bounds for the jackknife Wald interval, extending the Berry–Esseen approach of Li (2026) to the jackknife pivotal quantity. Second, a formal comparison of coverage accuracy between the BCa bootstrap and the jackknife under Edgeworth expansions for the ALE pivot. Third, an analytic estimator for $c_{R}/\sigma^{2}_{\mathrm{EIF}}$ that avoids full jackknife re-fitting and provides a faster diagnostic.
Conclusion.
The near-boundary regime is not an exotic edge case. It is the default operating regime for semiparametric estimators at practically relevant sample sizes and nuisance library configurations. A variance ratio of 1.5–2.0 between total and first-order variance, corresponding to confidence interval undercoverage of 4–8 percentage points, is empirically common and theoretically expected at the product-rate boundary. The refined procedures developed here provide a principled, computationally tractable, and asymptotically grounded resolution.
Appendix A Proof of Proposition 6.1 (ICC Amplification)
We sketch the argument for the SA-TMLE DGP. Write the remainder as:
$R_{2}=\frac{1}{J}\sum_{j=1}^{J}\tilde{R}_{j},\quad\tilde{R}_{j}=\sum_{i,t}\bigl(\hat{\bar{Q}}_{Y}(S_{ijt},a,W_{ijt},t)-\bar{Q}_{Y}^{0}(\cdot)\bigr)\Bigl(\tfrac{\hat{g}_{\Delta}(\cdot)}{{g^{0}_{\Delta}}(\cdot)}-1\Bigr).$
The variance of $R_{2}$ is:
$\mathrm{Var}(R_{2})=\frac{1}{J^{2}}\sum_{j,j^{\prime}}\mathrm{Cov}(\tilde{R}_{j},\tilde{R}_{j^{\prime}})=\frac{1}{J}\mathrm{Var}(\tilde{R}_{j}),$
using cluster independence for $j\neq j^{\prime}$. Within cluster $j$, the observations share $b_{j}$, inducing correlation between the nuisance residuals $(\hat{\bar{Q}}_{Y}-\bar{Q}_{Y}^{0})(S_{ijt},\cdot)$ and $(\hat{\bar{Q}}_{Y}-\bar{Q}_{Y}^{0})(S_{ij^{\prime}t^{\prime}},\cdot)$ for $(i,t)\neq(i^{\prime},t^{\prime})$ in the same cluster. This between-observation correlation within a cluster contributes a term proportional to $\rho_{\mathrm{ICC}}\times m\times(\mathrm{average\;pairwise\;covariance\;of\;nuisance\;residuals})$, establishing (30).