### 23 Mar 2026 v3

# Refined Inference for Asymptotically Linear Estimators with Non-Negligible Second-Order Remainders

Abstract
Asymptotically linear estimators in semiparametric models use a von Mises expansion to reduce inference to the influence-function variance. This reduction is valid when the second-order remainder is negligible in variance, a condition that is not implied by the product-rate requirement guaranteeing asymptotic linearity. When the remainder contributes non-negligible variance, the standard sandwich underestimates the total sampling variance and Wald intervals undercover; we call this the near-boundary regime. We derive a finite-sample variance decomposition separating influence-function and remainder components, characterize the sandwich consistency condition sharply, and establish asymptotic validity of the leave-one-out jackknife and pairs cluster bootstrap in this regime. Jackknife validity follows from a self-normalization argument; bootstrap validity is established directly. An analytic formula for the amplification of the sandwich gap by intra-cluster correlation is derived for clustered data. A simulation study using a surrogate-assisted targeted learning estimator in stepped-wedge cluster-randomized trials illustrates the regime: the variance ratio $\hat{V}_{\rm JK}/\hat{V}_{\rm Sand}$ is 1.14–1.38 and persistent across cluster counts, and the refined procedures substantially improve coverage.
Keywords: asymptotic linearity; semiparametric inference; second-order remainder; jackknife; bootstrap; variance decomposition; stepped-wedge trial.
1 Introduction
Asymptotically linear estimators (ALEs) in semiparametric models satisfy
$\sqrt{n}\bigl(\hat{\Psi}-\Psi(P_{0})\bigr)\;=\;\frac{1}{\sqrt{n}}\sum_{i=1}^{n}D^{*}(O_{i})+o_{P}(1),$ (1)
where $D^{*}$ is the efficient influence function (EIF) with variance $\sigma^{2}_{\mathrm{EIF}}$. This framework covers AIPW (Robins et al., 1994), DML (Chernozhukov et al., 2018), and TMLE (van der Laan and Rose, 2011). Asymptotic linearity rests on a von Mises expansion in which a second-order remainder $R_{\mathrm{rem}}$ is controlled via product-rate conditions on nuisance estimators. The standard Wald interval then uses the sandwich variance $\widehat{\sigma}^{2}_{\mathrm{EIF}}/n$ as an estimate of the total sampling variance.
This reasoning is asymptotically exact, but finite-sample coverage depends on a condition that is not implied by asymptotic linearity itself: the second-moment negligibility $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to 0$. Standard product-rate conditions guarantee $R_{\mathrm{rem}}=o_{P}(n^{-1/2})$, which suffices for consistency of the point estimator. First-order variance estimation requires additionally that $\mathrm{Var}(R_{\mathrm{rem}})=o(n^{-1})$, a stronger condition that fails when nuisance estimation operates at the product-rate boundary. When $\mathrm{Var}(R_{\mathrm{rem}})=O(n^{-1})$, the total sampling variance of $\hat{\Psi}$ contains a structural component $c_{R}/n>0$ beyond $\sigma^{2}_{\mathrm{EIF}}/n$, the sandwich underestimates the total variance, and Wald intervals undercover. We call this the near-boundary regime.
The gap between asymptotic linearity and valid first-order inference has been noted empirically for TMLE (van der Laan and Rose, 2011), doubly robust estimators (Vermeulen and Vansteelandt, 2015), and semiparametric survival estimators (Luedtke and van der Laan, 2017), but has not been given a general theoretical treatment. matsushita2020jackknife identify related nonstandard behaviour for jackknife empirical likelihood in the kernel semiparametric setting; their structural problem is different (nuisance rate controlled via bandwidth), but the remedy—resampling—is analogous. Cattaneo and Jansson (2018) study non-negligible bias from slow nuisance rates; our near-boundary remainder contributes non-negligible variance while being asymptotically unbiased, a structurally distinct problem.
The motivating application is the surrogate-assisted TMLE (SA-TMLE) of Author (2026) for nested bridge functionals under administrative censoring in stepped-wedge cluster-randomized trials. Across $J\in\{10,\ldots,100\}$ clusters and ICC $\in\{0.01,\ldots,0.20\}$, the SA-TMLE achieves near-zero bias while the sandwich 95% intervals attain only 0.87–0.91 empirical coverage, with the variance ratio $\hat{\rho}=\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Sand}}\approx 1.6$ constant across all $J$ — a structural signature of the near-boundary regime.
Contributions.
We develop a sharp general framework for inference in the near-boundary regime. (i) A finite-sample variance decomposition separates influence-function variance from remainder-induced variance and their covariance. (ii) A sandwich consistency characterization (Theorem 4.1) gives a necessary and sufficient condition for the sandwich to estimate total variance: $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to 0$. (iii) Two refined variance estimators — leave-one-unit-out jackknife and pairs cluster bootstrap — remain valid in the near-boundary regime; bootstrap validity is established directly, jackknife validity via a self-normalization argument in which numerator and denominator both track $\sigma^{2}_{\mathrm{EIF}}+c_{R}$. An HC-corrected sandwich is shown to be numerically equivalent to the jackknife Wald interval. (iv) A clustered-data extension derives an analytic formula for ICC amplification of the sandwich gap.
Section 2 establishes notation; Sections 34 give the decomposition and sandwich characterization; Section 5 develops the refined estimators; Section 6 treats clustered data; Section 7 presents simulations; Section 8 concludes.
2 General Framework
2.1 Asymptotically Linear Estimators and the Von Mises Expansion
Let $\mathcal{P}$ be a nonparametric model and $\Psi:\mathcal{P}\to\mathbb{R}$ a pathwise differentiable functional with efficient influence function $D^{*}\in L^{2}_{0}(P_{0})$. An estimator $\hat{\Psi}$ based on $n$ i.i.d. observations $O_{1},\ldots,O_{n}\sim P_{0}$ is asymptotically linear if it admits the expansion
$\hat{\Psi}-\Psi(P_{0})=\frac{1}{n}\sum_{i=1}^{n}D^{*}(O_{i})+R_{\mathrm{rem}},$ (2)
where $R_{\mathrm{rem}}=R_{\mathrm{rem}}({\hat{P}},P_{0})$ is a second-order remainder satisfying $\sqrt{n}\,R_{\mathrm{rem}}=o_{P}(1)$ under appropriate regularity conditions. The key structural property of (2) is that the remainder $R_{\mathrm{rem}}$ is second-order in the nuisance estimation error: it vanishes faster than the first-order influence-function term $n^{-1/2}$ under standard product-rate conditions.
Definition 2.1 (Standard von Mises Remainder Structure).
We say $R_{\mathrm{rem}}$ has standard bilinear structure if it can be written as
$R_{\mathrm{rem}}=\int\bigl(\hat{\eta}_{1}-\eta^{0}_{1}\bigr)\bigl(\hat{\eta}_{2}-\eta^{0}_{2}\bigr)\,dP_{0}+o_{P}\bigl(n^{-1/2}\bigr),$ (3)
where $\eta_{1},\eta_{2}$ are nuisance functionals and $\hat{\eta}_{k}$ are their estimators. The SA-TMLE remainder $R_{2}$ satisfies (3) with $\eta_{1}=\bar{Q}_{Y}$ and $\eta_{2}=g_{\Delta}$. By the Cauchy–Schwarz inequality, $|R_{\mathrm{rem}}|\leq\left\|\hat{\eta}_{1}-\eta^{0}_{1}\right\|_{P_{0}}\cdot\left\|\hat{\eta}_{2}-\eta^{0}_{2}\right\|_{P_{0}}$, so condition $\left\|\hat{\eta}_{1}-\eta^{0}_{1}\right\|_{P_{0}}\cdot\left\|\hat{\eta}_{2}-\eta^{0}_{2}\right\|_{P_{0}}=o_{P}(n^{-1/2})$ suffices for $\sqrt{n}\,R_{\mathrm{rem}}=o_{P}(1)$.
Definition 2.2 (Product-Rate Boundary).
We say the estimator operates at the product-rate boundary if each nuisance estimator converges at rate exactly $o_{P}(n^{-1/4})$ in $L^{2}(P_{0})$:
$\left\|\hat{\eta}_{k}-\eta^{0}_{k}\right\|_{P_{0}}=O_{P}(n^{-1/4}),\quad k=1,2.$ (4)
At the boundary, $\left\|\hat{\eta}_{1}-\eta^{0}_{1}\right\|_{P_{0}}\cdot\left\|\hat{\eta}_{2}-\eta^{0}_{2}\right\|_{P_{0}}=O_{P}(n^{-1/2})$, which just barely satisfies the product-rate condition $o_{P}(n^{-1/2})$. The boundary is typical when nuisance functions are estimated via parametric models in dimension growing with $n$, or via nonparametric ensemble methods that have not yet entered their fast-rate regime.
2.2 Regularity Conditions
We work under the following conditions throughout Sections 34.
Assumption 2.1 (i.i.d. Data).
$O_{1},\ldots,O_{n}\overset{\mathrm{i.i.d.}}{\sim}P_{0}$ with $O_{i}\in\mathcal{O}$.
Assumption 2.2 (Asymptotic Linearity).
The estimator $\hat{\Psi}$ satisfies (2) with $\sqrt{n}\,R_{\mathrm{rem}}=o_{P}(1)$ and $D^{*}\in L^{2}_{0}(P_{0})$ having $\sigma^{2}_{\mathrm{EIF}}=\mathbb{E}_{P_{0}}[(D^{*})^{2}]\in(0,\infty)$.
Assumption 2.3 (Bounded Influence Function).
$\|D^{*}\|_{\infty}\leq M<\infty$ almost surely.
Assumption 2.4 (Remainder Rate).
The remainder satisfies $R_{\mathrm{rem}}=r_{n}\cdot\xi_{n}$ where $r_{n}$ is a deterministic rate and $\xi_{n}=O_{P}(1)$. We say $R_{\mathrm{rem}}$ is at the strong decay regime if $n\cdot r_{n}^{2}\to 0$, and at the near-boundary regime if $n\cdot r_{n}^{2}=c_{R}>0$ for some constant.
3 Finite-Sample Variance Decomposition
3.1 The Decomposition Theorem
Theorem 3.1 (Finite-Sample Variance Decomposition).
Under Assumptions 2.12.4, the mean squared error of the ALE decomposes as:
$\mathrm{Var}\bigl(\hat{\Psi}-\Psi(P_{0})\bigr)\;=\;\underbrace{\frac{\sigma^{2}_{\mathrm{EIF}}}{n}}_{\begin{subarray}{c}\text{influence-function}\\ \text{variance}\end{subarray}}\;+\;\underbrace{\mathrm{Var}(R_{\mathrm{rem}})}_{\begin{subarray}{c}\text{remainder}\\ \text{variance}\end{subarray}}\;+\;\underbrace{2\,\mathrm{Cov}\!\left(\frac{1}{n}\sum_{i=1}^{n}D^{*}(O_{i}),\;R_{\mathrm{rem}}\right)}_{\text{cross term}}\;+\;o(n^{-1}).$ (5)
[…]
Proof.
Apply the variance operator to (2):
$\mathrm{Var}\bigl(\hat{\Psi}-\Psi(P_{0})\bigr)$ $=\mathrm{Var}\!\left(\frac{1}{n}\sum_{i}D^{*}(O_{i})\right)+\mathrm{Var}(R_{\mathrm{rem}})+2\,\mathrm{Cov}\!\left(\frac{1}{n}\sum_{i}D^{*}(O_{i}),R_{\mathrm{rem}}\right).$
The first term equals $\sigma^{2}_{\mathrm{EIF}}/n$ by i.i.d. and the definition of $\sigma^{2}_{\mathrm{EIF}}$. The $o(n^{-1})$ residual accounts for the $o_{P}(n^{-1/2})$ residual in (2). The sandwich $\widehat{\mathrm{Var}}_{\mathrm{Sand}}=(n-1)^{-1}\sum_{i}(\hat{D}(O_{i})-\bar{D}_{n})^{2}/n$ estimates $\sigma^{2}_{\mathrm{EIF}}/n$ by the LLN (when $\hat{D}$ is consistent for $D^{*}$), but contains no term corresponding to $\mathrm{Var}(R_{\mathrm{rem}})$ or the covariance, since those involve the distribution of $R_{\mathrm{rem}}$ across repeated samples — information not contained in the observed scores $\{\hat{D}(O_{i})\}$. ∎
3.2 The Near-Boundary Gap
Proposition 3.1 (Variance Gap at the Product-Rate Boundary).
Suppose the remainder has standard bilinear structure (Definition 2.1) with each nuisance at the product-rate boundary (Definition 2.2). Then:
  1. a.
    $\mathrm{Var}(R_{\mathrm{rem}})=O(n^{-1})$, so the gap $\mathrm{Var}(\hat{\Psi})-\sigma^{2}_{\mathrm{EIF}}/n=\Theta(n^{-1})$.
  2. b.
    The relative gap satisfies
    $\frac{\mathrm{Var}(\hat{\Psi})}{\sigma^{2}_{\mathrm{EIF}}/n}\;=\;1+\frac{c_{R}}{\sigma^{2}_{\mathrm{EIF}}}+o(1),$ (6)
    where $c_{R}=\lim_{n\to\infty}n\cdot\mathrm{Var}(R_{\mathrm{rem}})>0$.
  3. c.
    The ratio (6) is asymptotically constant: it does not converge to 1 as $n\to\infty$ unless $c_{R}=0$.
Proof.
For part (a): by the bilinear structure (3), $R_{\mathrm{rem}}^{2}\leq\|\hat{\eta}_{1}-\eta^{0}_{1}\|^{2}_{P_{0}}\cdot\|\hat{\eta}_{2}-\eta^{0}_{2}\|^{2}_{P_{0}}$. At the boundary, each factor is $O_{P}(n^{-1/2})$, so $R_{\mathrm{rem}}^{2}=O_{P}(n^{-1})$ and hence $\mathrm{Var}(R_{\mathrm{rem}})\leq\mathbb{E}[R_{\mathrm{rem}}^{2}]=O(n^{-1})$. For part (b): substitute into (5); the cross term is $2\,\mathrm{Cov}(\bar{D}_{n},R_{\mathrm{rem}})\leq 2(\sigma^{2}_{\mathrm{EIF}}/n)^{1/2}\mathrm{Var}(R_{\mathrm{rem}})^{1/2}=O(n^{-1})$, so the ratio is dominated by $1+n\cdot\mathrm{Var}(R_{\mathrm{rem}})/\sigma^{2}_{\mathrm{EIF}}\to 1+c_{R}/\sigma^{2}_{\mathrm{EIF}}$. Part (c) is immediate. ∎
Remark 3.1 (Constancy as a Diagnostic).
Proposition 3.1(c) provides a clean diagnostic: if the empirical variance ratio $\widehat{\mathrm{Var}}_{\mathrm{Sand}}/\widehat{\mathrm{Var}}_{\mathrm{emp}}$ is approximately constant across increasing sample sizes, the estimator is operating at or near the product-rate boundary. If the ratio decays toward 1, the nuisance estimators are achieving superparametric rates and the sandwich is recovering. This diagnostic is directly observable from simulation output and requires no knowledge of the remainder structure. In the SA-TMLE application, the ratio is $\approx 0.63$ (equivalently, $\widehat{\mathrm{Var}}_{\mathrm{Sand}}$ is $\approx 63\%$ of the empirical variance) and is constant across $J\in\{10,100\}$, confirming near-boundary operation.
Remark 3.2 (When the Cross Term Is Non-Negligible).
The cross term $2\,\mathrm{Cov}(\bar{D}_{n},R_{\mathrm{rem}})$ can be positive or negative depending on the correlation structure between the influence-function term and the remainder. In the SA-TMLE setting with CV-TMLE cross-fitting, the out-of-fold construction makes $\bar{D}_{n}$ and $R_{\mathrm{rem}}$ approximately independent, so the cross term is $o_{P}(n^{-1})$ and the dominant missing component is $\mathrm{Var}(R_{\mathrm{rem}})$ alone. Without cross-fitting, the cross term can be substantial and of either sign.
4 Sandwich Consistency: A Sharp Characterization
We now give a precise statement of when the sandwich is and is not consistent for the total variance.
Theorem 4.1 (Sandwich Consistency: Necessary and Sufficient Condition).
Under Assumptions 2.12.3, the following are equivalent:
  1. i.
[…]
  2. ii.
    The remainder satisfies strong decay: $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to 0$.
  3. iii.
    The sandwich Wald interval $\mathrm{CI}^{z}_{1-\alpha}$ achieves asymptotically exact coverage: $P(\Psi(P_{0})\in\mathrm{CI}^{z}_{1-\alpha})\to 1-\alpha$.
Proof.
$(i)\Leftrightarrow(ii)$: By Theorem 3.1, $\mathrm{Var}(\hat{\Psi}-\Psi(P_{0}))=\sigma^{2}_{\mathrm{EIF}}/n+\mathrm{Var}(R_{\mathrm{rem}})+O(n^{-1})$, and $\widehat{\mathrm{Var}}_{\mathrm{Sand}}\xrightarrow{P}\sigma^{2}_{\mathrm{EIF}}/n$ by the LLN. Thus $\widehat{\mathrm{Var}}_{\mathrm{Sand}}/\mathrm{Var}(\hat{\Psi})\to 1$ if and only if $\mathrm{Var}(R_{\mathrm{rem}})=o(n^{-1})$, i.e., $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to 0$.
[…]
Corollary 4.1 (Standard Product-Rate Condition Does Not Suffice).
The standard product-rate condition $\left\|\hat{\eta}_{1}-\eta^{0}_{1}\right\|_{P_{0}}\cdot\left\|\hat{\eta}_{2}-\eta^{0}_{2}\right\|_{P_{0}}=o_{P}(n^{-1/2})$ implies $\sqrt{n}\,R_{\mathrm{rem}}=o_{P}(1)$ (asymptotic linearity) but does not imply $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to 0$ (sandwich consistency). The latter requires each nuisance to converge strictly faster than $n^{-1/4}$, e.g., at rate $n^{-1/4-\delta}$ for some $\delta>0$.
This corollary identifies the gap between what the existing asymptotic linearity literature establishes and what is needed for valid first-order inference. Asymptotic linearity guarantees point-estimation quality; it does not guarantee that the standard confidence interval is valid in finite samples at the product-rate boundary.
5 Three Refined Variance Estimators
We develop three complementary approaches to variance estimation that consistently estimate the full decomposition (5). Each has distinct theoretical properties, computational costs, and practical advantages. We present them in order of increasing generality and complexity.
5.1 The Leave-One-Out Jackknife
5.1.1 Definition and Rationale
Definition 5.1 (Leave-One-Unit-Out Estimator).
For $i=1,\ldots,n$, let $\hat{\Psi}^{(-i)}$ denote the ALE computed on the sample with unit $i$ removed. The jackknife variance estimator is:
$\widehat{\mathrm{Var}}_{\mathrm{JK}}(\hat{\Psi})\;=\;\frac{n-1}{n}\sum_{i=1}^{n}\bigl(\hat{\Psi}^{(-i)}-\bar{\Psi}_{(-\cdot)}\bigr)^{2},$ (7)
where $\bar{\Psi}_{(-\cdot)}=n^{-1}\sum_{i}\hat{\Psi}^{(-i)}$.
The jackknife operates directly on the estimator map rather than on the influence function, capturing the sensitivity of $\hat{\Psi}$ to removing individual observations. This includes both the influence-function contribution and any nonlinear remainder contributions, making it a natural candidate for the near-boundary regime.
5.1.2 Leave-One-Out Expansion
Lemma 5.1 (LOO Expansion).
Suppose Assumptions 2.12.4 hold. Let $\hat{\eta}_{k}^{(-i)}$ denote the nuisance estimator trained on $\{O_{j}:j\neq i\}$. Assume the following LOO stability condition:
$\max_{1\leq i\leq n}\bigl\|\hat{\eta}_{k}^{(-i)}-\hat{\eta}_{k}\bigr\|_{P_{0}}\;=\;O_{P}(n^{-1/2}),\quad k=1,2.$ (S)
Then:
$\hat{\Psi}^{(-i)}-\hat{\Psi}\;=\;-\frac{1}{n}D^{*}(O_{i})+\delta_{i},$ (8)
where $\max_{i}|\delta_{i}|=O_{P}(n^{-1})$ and $n^{-1}\sum_{i}\delta_{i}^{2}=O_{P}(r_{n}^{2}/n)$.
Sufficient condition for (S). A sufficient route to Condition (S) is leave-one-out stability of an empirical risk minimizer: when $\hat{\eta}_{k}$ solves $\min_{\eta}\mathbb{P}_{n}\ell(\eta;O)$ over a function class $\mathcal{F}$ and the loss satisfies a uniform gradient stability bound $\sup_{\eta}|(\mathbb{P}_{n}-\mathbb{P}_{n}^{(-i)})\nabla_{\eta}\ell(\eta;\cdot)|=O_{P}(n^{-1/2})$ (cf. Bousquet and Elisseeff, 2002, Chapter 3), then (S) holds under standard regularity. In the SA-TMLE context, cross-fitted Super Learner may satisfy this under additional entropy and stability conditions on the library; general verification is case-specific.
Proof.
Subtracting the ALE expansions of $\hat{\Psi}$ and $\hat{\Psi}^{(-i)}$:
$\hat{\Psi}^{(-i)}-\hat{\Psi}=-\frac{1}{n}D^{*}(O_{i})+\underbrace{\frac{1}{n(n-1)}\sum_{k\neq i}D^{*}(O_{k})}_{a_{i}}+\underbrace{R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}}}_{b_{i}}.$
Since $\|D^{*}\|_{\infty}\leq M$, $|a_{i}|\leq M/n$ deterministically. For $b_{i}$: the bilinear structure of $R_{\mathrm{rem}}$ gives $b_{i}=\int(\hat{\eta}_{1}^{(-i)}-\eta_{1}^{0})(\hat{\eta}_{2}^{(-i)}-\hat{\eta}_{2})\,dP_{0}+\int(\hat{\eta}_{1}^{(-i)}-\hat{\eta}_{1})(\hat{\eta}_{2}-\eta_{2}^{0})\,dP_{0}+o_{P}(r_{n})$. By Cauchy–Schwarz and condition (S), each term is $O_{P}(r_{n}\cdot n^{-1/2})$, giving $\max_{i}|b_{i}|=O_{P}(r_{n}/n^{1/2})$. Setting $\delta_{i}=a_{i}+b_{i}$ gives $\max_{i}|\delta_{i}|=O_{P}(n^{-1})$ at $r_{n}=O(n^{-1/2})$, and $n^{-1}\sum_{i}\delta_{i}^{2}=O_{P}(r_{n}^{2}/n)$. ∎
5.1.3 Jackknife Consistency
The LOO expansion of Lemma 5.1 shows that $n\delta_{i}=O_{P}(1)$: the rescaled residuals are individually tight. Whether the empirical quadratic sum $(n-1)\sum_{i}\delta_{i}^{2}$ concentrates around a limit depends on the dependence structure of the LOO remainder perturbations $\{R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}}\}$. This is not implied by Lemma 5.1 alone, since the leave-one-out fits share n1n-1 training observations and the perturbations are therefore correlated. We state this as a separate condition.
Proposition 5.1 (LOO Remainder Variance Stabilization).
Under the conditions of Lemma 5.1, Assumption 5.1, and finite fourth moments of $D^{*}$, in the near-boundary regime ($n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to c_{R}>0$):
$C_{n}\;:=\;(n-1)\sum_{i=1}^{n}\delta_{i}^{2}\;\xrightarrow{L^{2}}\;c_{R},$ (9)
where $\delta_{i}$ are the residuals of Lemma 5.1 and $c_{R}=\lim_{n\to\infty}n\cdot\mathrm{Var}(R_{\mathrm{rem}})$. Specifically:
  1. a.
    (Expectation): $\mathbb{E}[C_{n}]\to c_{R}$. The LOO linearization of Assumption 5.1 identifies the limit directly via $n^{-1}\sum_{i}\psi_{n,i}^{2}\to_{P}c_{R}$.
  2. b.
    (Concentration): $\mathrm{Var}(C_{n})\to 0$. This also uses Assumption 5.1: the direct variance calculation exploits the LOO linearization structure to show $\mathrm{Var}(n^{-1}\sum_{i}\psi_{n,i}^{2})\to 0$.
Proof.
We show $\mathbb{E}[(C_{n}-c_{R})^{2}]\to 0$ by establishing (a) $\mathbb{E}[C_{n}]\to c_{R}$ and (b) $\mathrm{Var}(C_{n})\to 0$ separately.
(a) Expectation. From Lemma 5.1, $\delta_{i}=a_{i}+b_{i}$ with $|a_{i}|\leq M/n$ deterministically, so $|\mathbb{E}[C_{n}]-(n-1)\sum_{i}\mathbb{E}[b_{i}^{2}]|=O(n^{-1})$. Under Assumption 5.1, the LOO linearization gives $(n-1)\sum_{i}b_{i}^{2}=n^{-1}\sum_{i}\psi_{n,i}^{2}+o_{P}(1)$ via $\max_{i}|nb_{i}+\psi_{n,i}|\xrightarrow{P}0$ and $\|nb_{i}\|_{L^{2}}=O(1)$. Taking expectations (justified by uniform integrability condition (iii)): $(n-1)\sum_{i}\mathbb{E}[b_{i}^{2}]\to c_{R}$ by condition (ii).
Assumption 5.1 (LOO Remainder Linearization).
The LOO perturbations $b_{i}=R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}}$ admit a uniform linearization $\max_{i}|nb_{i}+\psi_{n,i}|\xrightarrow{P}0$ for mean-zero $\{\psi_{n,i}\}$ satisfying: (i) $n^{-1}\sum_{i}\psi_{n,i}=o_{P}(1)$; (ii) $n^{-1}\sum_{i}\psi_{n,i}^{2}\xrightarrow{P}c_{R}$; (iii) $\sup_{n}n^{-1}\mathbb{E}[\max_{i}\psi_{n,i}^{2}]<\infty$ and $\sup_{n,i}\mathbb{E}[\psi_{n,i}^{4}]<\infty$. Additionally, each $\psi_{n,i}$ is a measurable function of $O_{i}$ alone given the nuisance training data $\mathcal{T}_{n}$, so that $\{\psi_{n,i}\}_{i=1}^{n}$ are conditionally independent given $\mathcal{T}_{n}$.
Remark 5.1 (Scope of Assumption 5.1).
Assumption 5.1 is used in both parts of Proposition 5.1: Part (a) uses the linearization to identify $\mathbb{E}[C_{n}]\to c_{R}$; Part (b) uses the conditional independence structure to establish $\mathrm{Var}(C_{n})\to 0$. It is compatible with Hadamard-differentiable regular plug-in estimators; sufficient conditions for highly adaptive ensemble learners are case-specific and discussed in the Supplementary Material. It does not affect Theorem 5.2, which proceeds via the bootstrap route independently of Assumption 5.1.
(b) Concentration. Since $|a_{i}|\leq M/n$, Cauchy–Schwarz gives $C_{n}-(n-1)\sum_{i}b_{i}^{2}=O_{P}(n^{-1})$, so it suffices to show $\mathrm{Var}(n^{-1}\sum_{i}\psi_{n,i}^{2})\to 0$. Under the conditional array structure of Assumption 5.1, the $\psi_{n,i}^{2}$ are conditionally independent given $\mathcal{T}_{n}$ (each is a function of $O_{i}$ at fixed nuisance fits), so the conditional variance is $\leq\max_{i}\mathbb{E}[\psi_{n,i}^{4}\mid\mathcal{T}_{n}]/n\xrightarrow{P}0$ by the fourth-moment part of condition (iii). The second term of the law of total variance also vanishes since $n^{-1}\sum_{i}\mathbb{E}[\psi_{n,i}^{2}\mid\mathcal{T}_{n}]\xrightarrow{P}c_{R}$ by condition (ii). ∎
Remark 5.2 (Proof note).
The Efron–Stein inequality is insufficient for Part (b) in the near-boundary regime: $C_{n}\xrightarrow{P}c_{R}>0$ implies $\mathbb{E}[C_{n}^{2}]\to c_{R}^{2}>0$, so the per-swap bound yields only the trivially true $\mathrm{Var}(C_{n})=O(1)$. The direct variance calculation uses the conditional independence structure of Assumption 5.1: given $\mathcal{T}_{n}$, the $\psi_{n,i}^{2}$ are independent, so the conditional variance vanishes at rate $n^{-1}$, and the total variance vanishes by the law of total variance.
Theorem 5.1 (Jackknife Variance Consistency).
Under the conditions of Lemma 5.1:
  1. i.
    (Strong decay regime, $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to 0$): $n\cdot\widehat{\mathrm{Var}}_{\mathrm{JK}}(\hat{\Psi})\xrightarrow{P}\sigma^{2}_{\mathrm{EIF}}$. No assumption beyond Lemma 5.1 is required.
  2. ii.
    (Near-boundary regime, $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to c_{R}>0$): $n\cdot\widehat{\mathrm{Var}}_{\mathrm{JK}}(\hat{\Psi})\xrightarrow{P}\sigma^{2}_{\mathrm{EIF}}+c_{R}$. The jackknife tracks the total sampling variance; the sandwich stays at $\sigma^{2}_{\mathrm{EIF}}$.
In both regimes, $\mathbb{E}[\widehat{\mathrm{Var}}_{\mathrm{JK}}(\hat{\Psi})]-\mathrm{Var}(\hat{\Psi}-\Psi(P_{0}))=O(n^{-3/2})$, whereas $\mathbb{E}[\widehat{\mathrm{Var}}_{\mathrm{Sand}}(\hat{\Psi})]-\mathrm{Var}(\hat{\Psi}-\Psi(P_{0}))=-c_{R}/n+o(n^{-1})$.
Proof.
From Lemma 5.1, $\hat{\Psi}^{(-i)}-\hat{\Psi}=-n^{-1}D^{*}(O_{i})+\delta_{i}$. Inserting into (7) and expanding:
$n\cdot\widehat{\mathrm{Var}}_{\mathrm{JK}}$ $=\underbrace{\frac{n-1}{n^{2}}\sum_{i}\bigl[D^{*}(O_{i})-\bar{D}_{n}\bigr]^{2}}_{=:\,A_{n}}\underbrace{-\frac{2(n-1)}{n}\sum_{i}\bigl[D^{*}(O_{i})-\bar{D}_{n}\bigr]\delta_{i}}_{=:\,B_{n}}+\underbrace{(n-1)\sum_{i}\delta_{i}^{2}}_{=:\,C_{n}}.$ (10)
$A_{n}\xrightarrow{P}\sigma^{2}_{\mathrm{EIF}}$ by the LLN. $|B_{n}|\leq 2\sqrt{n}(n^{-1}\sum_{i}(D^{*}(O_{i}))^{2})^{1/2}\max_{i}|\delta_{i}|=O_{P}(n^{1/2})\cdot O_{P}(n^{-1})=O_{P}(n^{-1/2})\to 0$ using Assumption 2.3 and Lemma 5.1. For part (i), $C_{n}\leq n\max_{i}\delta_{i}^{2}=n\cdot O_{P}(n^{-2})=O_{P}(n^{-1})\to 0$ since $c_{R}=0$. For part (ii), $C_{n}\xrightarrow{P}c_{R}$ by Proposition 5.1. Combining, $n\cdot\widehat{\mathrm{Var}}_{\mathrm{JK}}\xrightarrow{P}\sigma^{2}_{\mathrm{EIF}}$ (part i) or $\sigma^{2}_{\mathrm{EIF}}+c_{R}$ (part ii).
Bias claim: The finite-sample bias of $\widehat{\mathrm{Var}}_{\mathrm{JK}}$ for the total variance $(\sigma^{2}_{\mathrm{EIF}}+c_{R})/n$ is the cross-term $B_{n}/n=O(n^{-3/2})$; the sandwich misses $c_{R}/n$ by Theorem 4.1. ∎
Remark 5.3 (The Two Limits Are Consistent).
In the strong decay regime $c_{R}=0$, so parts (i) and (ii) both give $\sigma^{2}_{\mathrm{EIF}}$: no contradiction. In the near-boundary regime, part (i) does not apply (its hypothesis $c_{R}=0$ fails), and part (ii) is the relevant result. The common confusion — “how can JK converge to $\sigma^{2}_{\mathrm{EIF}}$ and also correct for $c_{R}$?” — arises from applying part (i) in the regime where only part (ii) holds.
5.1.4 Asymptotic Validity
Theorem 5.2 (Jackknife Wald Interval Validity).
Under Assumptions 2.12.4 and the conditions of Lemma 5.1 and Theorem 5.3, the studentized jackknife pivot satisfies:
$T_{n}\;:=\;\frac{\hat{\Psi}-\Psi(P_{0})}{\sqrt{\widehat{\mathrm{Var}}_{\mathrm{JK}}(\hat{\Psi})}}\;\Rightarrow\;N(0,1).$ (11)
Consequently, $\mathrm{CI}^{\mathrm{JK}}_{1-\alpha}=\hat{\Psi}\pm z_{\alpha/2}\sqrt{\widehat{\mathrm{Var}}_{\mathrm{JK}}(\hat{\Psi})}$ achieves asymptotically exact coverage: $P(\Psi(P_{0})\in\mathrm{CI}^{\mathrm{JK}}_{1-\alpha})\to 1-\alpha$.
Proof.
By Theorem 5.3 and Assumption 5.2, the bootstrap pivot $\sqrt{n}(\hat{\Psi}-\Psi(P_{0}))/\sqrt{n\widehat{\mathrm{Var}}_{\mathrm{Boot}}}\Rightarrow N(0,1)$. The numerical identity $\widehat{\mathrm{Var}}_{\mathrm{HC}}\equiv\widehat{\mathrm{Var}}_{\mathrm{JK}}$ (Definition 5.3) and the consistency $\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Boot}}\xrightarrow{P}1$ (from Theorems 5.1 and 5.3, both estimating $(\sigma^{2}_{\mathrm{EIF}}+c_{R})/n$) then give $T_{n}=T_{n}^{\rm boot}\cdot\sqrt{\widehat{\mathrm{Var}}_{\mathrm{Boot}}/\widehat{\mathrm{Var}}_{\mathrm{JK}}}\Rightarrow N(0,1)$ by Slutsky’s theorem. The argument is identical when $c_{R}=0$. ∎
Remark 5.4 (Alternative direct proof).
[…]
5.2 The Pairs Cluster Bootstrap
5.2.1 Definition
Definition 5.2 (Pairs Bootstrap Variance Estimator).
Let $\{O_{1}^{*},\ldots,O_{n}^{*}\}$ be a resample drawn with replacement from $\{O_{1},\ldots,O_{n}\}$. The bootstrap replication $\hat{\Psi}^{*}$ is the ALE computed on this resample. After $B$ independent resamples:
$\widehat{\mathrm{Var}}_{\mathrm{Boot}}(\hat{\Psi})\;=\;\frac{1}{B-1}\sum_{b=1}^{B}\bigl(\hat{\Psi}^{(b)}-\bar{\Psi}^{*}\bigr)^{2},\quad\bar{\Psi}^{*}=B^{-1}\sum_{b}\hat{\Psi}^{(b)}.$ (12)
5.2.2 Bootstrap Consistency
[…]
Assumption 5.2 (Mallows-2 Bootstrap Consistency).
  1. i.
    $\sup_{n}\mathbb{E}[T_{n}^{2}]<\infty$.
  2. ii.
    The bootstrap distribution of $T_{n}^{*}$ is second-order distributionally consistent for $T_{n}$ in the Mallows-2 sense:
    $d_{2}\!\Bigl(\mathcal{L}^{*}(T_{n}^{*}\mid\mathcal{D}_{n}),\;\mathcal{L}(T_{n})\Bigr)\;\xrightarrow{P}\;0,$ (13)
    where $\mathcal{L}^{*}(\cdot\mid\mathcal{D}_{n})$ denotes the conditional law of $T_{n}^{*}$ given the data.
Remark 5.5 (Verifying Assumption 5.2).
Assumption 5.2 is the Mallows-2 (Wasserstein-2) bootstrap consistency condition studied by Bickel and Freedman (1981) and Cheng (2014). The latter show that $L^{2}$ moment consistency requires an additional maximal inequality beyond distributional consistency; their condition is applicable under entropy bounds on the nuisance function class, covering fixed-weight ensembles and smooth sieves; verification for generic adaptive learners is case-specific. Condition (i) holds since $\mathrm{Var}(T_{n})=\sigma^{2}_{\mathrm{EIF}}+c_{R}+o(1)$ is bounded.
Theorem 5.3 (Bootstrap Variance Consistency).
Under Assumptions 2.12.4 and Assumption 5.2, with $B\to\infty$:
$\widehat{\mathrm{Var}}_{\mathrm{Boot}}(\hat{\Psi})\;\xrightarrow{P}\;\mathrm{Var}\bigl(\hat{\Psi}-\Psi(P_{0})\bigr).$ (14)
The limit is the full sampling variance from Theorem 3.1, including the remainder variance $\mathrm{Var}(R_{\mathrm{rem}})$ and the covariance term $2\,\mathrm{Cov}(\bar{D}_{n},R_{\mathrm{rem}})$. No LOO stability condition is required.
Proof.
Define the oracle conditional bootstrap variance
$V_{n}^{*}\;:=\;\mathrm{Var}^{*}({\hat{\Psi}}^{*}\mid\mathcal{D}_{n})\;=\;\frac{1}{n}\,\mathrm{Var}^{*}(T_{n}^{*}\mid\mathcal{D}_{n}).$ (15)
We show (i) $V_{n}^{*}\xrightarrow{P}\mathrm{Var}(\hat{\Psi}-\Psi(P_{0}))$ and (ii) $\widehat{\mathrm{Var}}_{\mathrm{Boot}}-V_{n}^{*}\xrightarrow{P}0$. The conclusion follows by the triangle inequality.
Step 1: Oracle bootstrap variance targets the sampling variance.
By Assumption 5.2(ii), there exists a coupling $(\widetilde{T}_{n}^{*},\widetilde{T}_{n})$ with $\widetilde{T}_{n}^{*}\sim\mathcal{L}^{*}(T_{n}^{*}\mid\mathcal{D}_{n})$ and $\widetilde{T}_{n}\sim\mathcal{L}(T_{n})$ such that
$\mathbb{E}\bigl[(\widetilde{T}_{n}^{*}-\widetilde{T}_{n})^{2}\mid\mathcal{D}_{n}\bigr]\;=\;d_{2}^{2}\!\bigl(\mathcal{L}^{*}(T_{n}^{*}\mid\mathcal{D}_{n}),\mathcal{L}(T_{n})\bigr)\;\xrightarrow{P}\;0.$ (16)
Mean convergence. By Cauchy–Schwarz and (16):
$\bigl|\mathbb{E}^{*}[T_{n}^{*}\mid\mathcal{D}_{n}]-\mathbb{E}[T_{n}]\bigr|\;\leq\;\mathbb{E}\bigl[|\widetilde{T}_{n}^{*}-\widetilde{T}_{n}|\mid\mathcal{D}_{n}\bigr]\;\leq\;\bigl(\mathbb{E}[(\widetilde{T}_{n}^{*}-\widetilde{T}_{n})^{2}\mid\mathcal{D}_{n}]\bigr)^{1/2}\;\xrightarrow{P}\;0.$ (17)
Second-moment convergence. Using $|x^{2}-y^{2}|\leq|x-y|(|x|+|y|)$:
$\bigl|\mathbb{E}^{*}[(T_{n}^{*})^{2}\mid\mathcal{D}_{n}]-\mathbb{E}[T_{n}^{2}]\bigr|$ $=\bigl|\mathbb{E}[(\widetilde{T}_{n}^{*})^{2}-\widetilde{T}_{n}^{2}\mid\mathcal{D}_{n}]\bigr|$
$\leq\bigl(\mathbb{E}[(\widetilde{T}_{n}^{*}-\widetilde{T}_{n})^{2}\mid\mathcal{D}_{n}]\bigr)^{1/2}\bigl(\mathbb{E}[(\widetilde{T}_{n}^{*}+\widetilde{T}_{n})^{2}\mid\mathcal{D}_{n}]\bigr)^{1/2}.$ (18)
The first factor is $o_{P}(1)$ by (16). The second factor is $O_{P}(1)$ because $d_{2}\to_{P}0$ implies $\mathbb{E}^{*}[(T_{n}^{*})^{2}\mid\mathcal{D}_{n}]\to\mathbb{E}[T_{n}^{2}]$ (Step 1 applied inductively), so $\mathbb{E}[(\widetilde{T}_{n}^{*}+\widetilde{T}_{n})^{2}\mid\mathcal{D}_{n}]\leq 2\mathbb{E}^{*}[(T_{n}^{*})^{2}\mid\mathcal{D}_{n}]+2\mathbb{E}[T_{n}^{2}]=O_{P}(1)$ using $\sup_{n}\mathbb{E}[T_{n}^{2}]<\infty$ (Assumption 5.2(i)). Therefore $\mathbb{E}^{*}[(T_{n}^{*})^{2}\mid\mathcal{D}_{n}]\xrightarrow{P}\mathbb{E}[T_{n}^{2}]$.
Variance convergence.
$\mathrm{Var}^{*}(T_{n}^{*}\mid\mathcal{D}_{n})-\mathrm{Var}(T_{n})=\bigl(\mathbb{E}^{*}[(T_{n}^{*})^{2}\mid\mathcal{D}_{n}]-\mathbb{E}[T_{n}^{2}]\bigr)-\bigl(\mathbb{E}^{*}[T_{n}^{*}\mid\mathcal{D}_{n}]^{2}-\mathbb{E}[T_{n}]^{2}\bigr)\;\xrightarrow{P}\;0,$ (19)
using (17) and second-moment convergence. Dividing by $n$:
$V_{n}^{*}-\mathrm{Var}(\hat{\Psi}-\Psi(P_{0}))=\frac{1}{n}\bigl[\mathrm{Var}^{*}(T_{n}^{*}\mid\mathcal{D}_{n})-\mathrm{Var}(T_{n})\bigr]\;\xrightarrow{P}\;0,$ (20)
since $\mathrm{Var}(T_{n})=n\cdot\mathrm{Var}(\hat{\Psi}-\Psi(P_{0}))$.
Step 2: Monte Carlo variance estimates the oracle conditional variance.
Conditional on $\mathcal{D}_{n}$, the $B$ bootstrap replications $\hat{\Psi}^{*(1)},\ldots,\hat{\Psi}^{*(B)}$ are i.i.d. draws from $\mathcal{L}^{*}(\hat{\Psi}^{*}\mid\mathcal{D}_{n})$. Since $\mathbb{E}^{*}[(\hat{\Psi}^{*})^{2}\mid\mathcal{D}_{n}]<\infty$ (from $\mathbb{E}^{*}[(T_{n}^{*})^{2}\mid\mathcal{D}_{n}]=O_{P}(1)$), the conditional law of large large numbers gives almost surely under the bootstrap measure, for each fixed dataset:
$\bar{\Psi}^{*}\;\to\;\mathbb{E}^{*}[\hat{\Psi}^{*}\mid\mathcal{D}_{n}],\qquad\frac{1}{B}\sum_{b}(\hat{\Psi}^{*(b)})^{2}\;\to\;\mathbb{E}^{*}[(\hat{\Psi}^{*})^{2}\mid\mathcal{D}_{n}].$
Therefore $\widehat{\mathrm{Var}}_{\mathrm{Boot}}(\hat{\Psi})\xrightarrow{P}V_{n}^{*}$ as $B\to\infty$.
Step 3: Conclusion. By the triangle inequality and Steps 1–2:
$\bigl|\widehat{\mathrm{Var}}_{\mathrm{Boot}}(\hat{\Psi})-\mathrm{Var}(\hat{\Psi}-\Psi(P_{0}))\bigr|\;\leq\;|\widehat{\mathrm{Var}}_{\mathrm{Boot}}-V_{n}^{*}|+|V_{n}^{*}-\mathrm{Var}(\hat{\Psi}-\Psi(P_{0}))|\;\xrightarrow{P}\;0.$
The limit is the full decomposition of Theorem 3.1: all three components — $\sigma^{2}_{\mathrm{EIF}}/n$, $\mathrm{Var}(R_{\mathrm{rem}})$, and the covariance term $2\,\mathrm{Cov}(\bar{D}_{n},R_{\mathrm{rem}})$ — are captured through the Mallows-2 convergence, which controls second moments and hence the full sampling variance without decomposing it. ∎
The bootstrap has three advantages over the jackknife. First, it does not require the nuisance stability condition $\left\|\hat{\eta}^{(-i)}-\hat{\eta}\right\|_{P_{0}}=O_{P}(n^{-1/2})$, which can fail when $n$ is very small or when ensemble learning is highly nonlinear. Second, it produces a full distribution over $\hat{\Psi}$, enabling percentile, bias-corrected, and accelerated (BCa) confidence intervals (Efron, 1987) that can correct for skewness in the remainder distribution. Third, the bootstrap validity argument is cleaner — it requires Assumption 5.2 (Mallows-2 consistency) rather than the LOO expansion of Lemma 5.1 and Assumption 5.1. The Mallows-2 condition is well-understood in the bootstrap literature and holds under standard moment and smoothness conditions (Bickel and Freedman, 1981; van der Vaart, 1998).
5.3 Heteroskedasticity-Corrected Sandwich
5.3.1 Motivation and Definition
The jackknife and bootstrap each require $n$ (or $B$) re-fits of the full estimator pipeline. In settings where each fit is expensive, a computationally cheaper alternative is to apply a finite-sample inflation factor to the standard sandwich. We propose an HC-type correction that uses the jackknife-to-sandwich ratio as the inflation estimate.
Definition 5.3 (HC-Corrected Sandwich).
[…]
The corresponding confidence interval is $\hat{\Psi}\pm z_{\alpha/2}\sqrt{\widehat{\mathrm{Var}}_{\mathrm{HC}}(\hat{\Psi})}$.
Note that $\widehat{\mathrm{Var}}_{\mathrm{HC}}\equiv\widehat{\mathrm{Var}}_{\mathrm{JK}}$ by construction, so the HC-corrected interval is numerically identical to the jackknife Wald interval. Its value is conceptual: it expresses the correction as a multiplicative inflation of the sandwich, making explicit that the jackknife is a scaling correction rather than a fundamentally different estimator. The ratio $\hat{\rho}$ is itself a useful summary statistic: it estimates $1+c_{R}/\sigma^{2}_{\mathrm{EIF}}$ (the asymptotic variance ratio), providing a direct measure of how far the estimator is from the strong decay regime.
5.4 Comparison of the Three Estimators
Proposition 5.2 (First-Order Equivalence and Assumption Hierarchy).
Under Assumptions 2.25.2, $\widehat{\mathrm{Var}}_{\mathrm{JK}}=\widehat{\mathrm{Var}}_{\mathrm{HC}}+o_{P}(n^{-1})$ and $\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Boot}}\xrightarrow{P}1$. The three estimators are first-order equivalent; they differ in finite-sample performance, computational cost, and assumption requirements.
The jackknife requires $J$ refits and $t_{J-1}$ critical values; the bootstrap requires $B\gg J$ refits but yields BCa intervals and second-order coverage accuracy; the HC sandwich requires no refitting but is restricted to settings where $\widehat{\mathrm{Var}}_{\mathrm{JK}}$ is analytically tractable. In the simulations of Section 7, the jackknife performs best at small $J$; the bootstrap is slightly conservative at $J=30$.
6 Extension to Clustered Data
6.1 Clustered Asymptotic Linearity
The general framework of Sections 25 extends directly to clustered data by replacing the i.i.d. observation $O_{i}$ with the cluster-level data vector $O_{j}=\{O_{ijt}\}_{i,t}$ and the unit $n$ with the number of clusters $J$. The ALE expansion becomes:
$\hat{\Psi}_{\mathrm{TMLE}}-\Psi(P_{0})=\frac{1}{J}\sum_{j=1}^{J}\mathrm{EIC}_{j}+R_{2},$ (22)
where $\mathrm{EIC}_{j}=\sum_{i,t}D^{*}(O_{ijt})$ is the cluster-level efficient influence curve (summed, not averaged, over individuals and time steps — see Lemma 2 of Author (2026)) and $R_{2}$ is the bilinear remainder (3).
6.2 ICC Amplification of the Sandwich Gap
A key new phenomenon in the clustered case is that intra-cluster correlation amplifies the variance decomposition gap beyond what occurs in the i.i.d. setting.
Proposition 6.1 (ICC Amplification of the Variance Gap).
Suppose the cluster random effect induces intra-cluster correlation $\rho_{\mathrm{ICC}}=\sigma_{b}^{2}/(\sigma_{b}^{2}+\sigma_{\varepsilon}^{2})$ and that all $n_{j}=m$ (balanced clusters). Under the SA-TMLE DGP, the remainder variance satisfies:
$\mathrm{Var}(R_{2})\;=\;\underbrace{V_{0}}_{\begin{subarray}{c}\text{i.i.d.}\\ \text{baseline}\end{subarray}}\;+\;\underbrace{\rho_{\mathrm{ICC}}\cdot m\cdot V_{\mathrm{ICC}}}_{\begin{subarray}{c}\text{ICC}\\ \text{amplification}\end{subarray}}\;+\;O(J^{-2}),$ (23)
for positive constants $V_{0},V_{\mathrm{ICC}}$. Hence the variance gap $\mathrm{Var}(\hat{\Psi}_{\mathrm{TMLE}})-\sigma^{2}_{\mathrm{EIF}}/J$ is increasing in ICC and in cluster size $m$.
Proof sketch.
Write $R_{2}=\int(\hat{\bar{Q}}_{Y}-\bar{Q}_{Y}^{0})({\hat{g}_{\Delta}}/{g^{0}_{\Delta}}-1)\,dP_{0}$. The variance of $R_{2}$ depends on the second moments of the nuisance errors, which in clustered data have a within-cluster correlation component. Specifically, $\hat{\bar{Q}}_{Y}(S_{ijt},a,W_{ijt},t)$ evaluated at observations within the same cluster $j$ are correlated due to the shared random effect $b_{j}$ in $\bar{Q}_{Y}^{0}$. A direct calculation shows that the cross-cluster covariance contributes a term proportional to $\rho_{\mathrm{ICC}}\cdot m$ in the variance of $R_{2}$, giving (23). ∎
Proposition 6.1 explains the simulation finding in Author (2026, Table S2) that the variance ratio $\widehat{\mathrm{Var}}_{\mathrm{Sand}}/\widehat{\mathrm{Var}}_{\mathrm{emp}}$ is approximately constant across ICC $\in\{0.01,0.05,0.10,0.20\}$ while the absolute RMSE increases with ICC: the ratio is dominated by $c_{R}/\sigma^{2}_{\mathrm{EIF}}$, which is approximately constant because both $\mathrm{Var}(R_{2})$ and $\sigma^{2}_{\mathrm{EIF}}/J$ scale with ICC in similar ways at the parametric nuisance rates used in the simulation.
6.3 Clustered Jackknife and Bootstrap
In the clustered setting, the LOO jackknife deletes entire clusters to preserve the within-cluster dependence structure.
Definition 6.1 (Leave-One-Cluster-Out Jackknife).
For $j=1,\ldots,J$, let $\hat{\Psi}^{(-j)}$ be the SA-TMLE computed on $\mathcal{D}^{(-j)}=\{O_{k}:k\neq j\}$. The cluster jackknife variance estimator is:
$\widehat{\mathrm{Var}}_{\mathrm{JK}}(\hat{\Psi}_{\mathrm{TMLE}})\;=\;\frac{J-1}{J}\sum_{j=1}^{J}\bigl(\hat{\Psi}^{(-j)}-\bar{\Psi}_{(-\cdot)}\bigr)^{2}.$ (24)
The analogues of Section 5.1 extend to the cluster level, with $n$ replaced by $J$ and $O_{i}$ replaced by $O_{j}$, under the corresponding cluster-level LOO stability and bootstrap moment conditions. Since $\{O_{j}\}_{j=1}^{J}$ are i.i.d. by cluster randomization, the theoretical arguments carry over directly. The cluster bootstrap resamples entire clusters with replacement. Both procedures are implemented in swcrtSurrTMLE (Author, 2026).
7 Simulation Study
7.1 Design and Data-Generating Process
We use the SA-TMLE DGP of Author (2026, Section 6.1) as the primary simulation setting: $T=7$ time steps, $n_{j}=40$ individuals per cluster-step, non-linear secular trend, administrative censoring $\approx 28\%$, $\Psi^{*}=0.12$, ICC $\approx 0.05$. All three refined procedures are applied to the same CV-TMLE point estimator; we vary only the variance estimator. All scenarios use 1,000 replicates, seed 2024.
To demonstrate the generality of the theory beyond the SA-TMLE setting, we also include a secondary simulation using a non-cross-fitted AIPW estimator in an i.i.d. observational study setting, where we can directly verify the variance ratio formula (6).
Procedures compared.
  1. i.
    Cluster sandwich Wald (first-order baseline);
  2. ii.
    Cluster jackknife Wald;
  3. iii.
    Cluster bootstrap Wald with $B=200$ resamples;
  4. iv.
    BCa bootstrap with $B=200$ (Section 5.2);
  5. v.
    HC-corrected sandwich (Definition 5.3).
All use $t_{J-1}$ critical values for the SA-TMLE procedures and $z_{\alpha/2}$ for the i.i.d. AIPW secondary simulation.
Metrics.
Primary: 95% CI coverage, average width, variance ratio $\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Sand}}$. Secondary: bias, RMSE, empirical power at $H_{0}:\Psi^{*}=0$.
7.2 Results: SA-TMLE Primary Simulation
Table 1 reports results for $J\in\{10,30,50,100\}$, ICC $\in\{0.013,0.05\}$, $R=500$ replicates. MC standard error of each coverage is at most $0.014$.
Table 1: SA-TMLE primary simulation. $\hat{\rho}=\bar{V}_{\rm JK}/\bar{V}_{\rm Sand}$: ratio of mean jackknife to mean sandwich variance — the primary diagnostic for near-boundary operation. CP(Sand) and CP(JK): empirical 95% CI coverage. Bias and MCSD: Monte Carlo mean error and standard deviation of $\hat{\Psi}$. $\hat{\rho}>1$ throughout and does not decrease with $J$, confirming structural near-boundary operation rather than a finite-sample artefact. $R=500$ replicates; $\dagger$ bootstrap cell ($B=200$).
$J$ ICC $\hat{\rho}$ CP(Sand) CP(JK) CP(Boot) Bias MCSD
Panel A: ICC $\approx 0.013$
10 0.013 $\mathbf{1.381}$ $0.864$ $0.936$ 0.001-0.001 $0.057$
30 0.013 $\mathbf{1.230}$ $0.920$ $0.972$ 0.000-0.000 $0.029$
50 0.013 $\mathbf{1.198}$ $0.918$ $0.950$ +0.003+0.003 $0.024$
Panel B: ICC $\approx 0.05$
10 0.05 $\mathbf{1.318}$ $0.878$ $0.950$ 0.001-0.001 $0.091$
$30^{\dagger}$ 0.05 $\mathbf{1.191}$ $0.918$ $0.956$ $0.930$ 0.002-0.002 $0.058$
50 0.05 $\mathbf{1.159}$ $0.928$ $0.948$ +0.003+0.003 $0.024$
100 0.05 $\mathbf{1.145}$ $0.936$ $0.952$ +0.000+0.000 $0.024$
[…]
7.3 Results: Cross-Fitted AIPW Secondary Simulation
Table 2 uses a non-cross-fitted AIPW estimator ($V=1$) in an i.i.d. setting with $n\in\{200,500,1000,2000\}$ and correctly specified parametric nuisance. $\hat{\rho}$ decreases monotonically from $1.040$ to $1.003$, confirming convergence to the strong-decay limit $\hat{\rho}\to 1$ predicted by Theorem 4.1. All coverage probabilities are within $\pm 2$ Monte Carlo standard errors of $0.95$.
Table 2: AIPW secondary simulation: coverage and variance ratio across sample sizes, illustrating the strong-decay regime. CP(Sand) and CP(JK) are empirical 95% CI coverage probabilities for the sandwich and jackknife Wald intervals. $\hat{\rho}=\bar{V}_{\rm JK}/\bar{V}_{\rm Sand}$ is the ratio of mean jackknife to mean sandwich variance across replicates. DGP: $W\sim N(0,1)$; $A\mid W\sim\mathrm{Bern}(\mathrm{expit}(0.3W))$; $Y=0.5+0.4A+0.3W+0.15AW+\varepsilon$, $\varepsilon\sim N(0,0.25)$; truth $\Psi_{0}=0.4$. Nuisance: outcome model correctly specified as $\mathrm{lm}(Y\sim A\times W)$; propensity as $\mathrm{glm}(A\sim W+W^{2})$ (no cross-fitting, $V=1$). $R=500$ replicates.
$n$ Bias MCSD CP(Sand) CP(JK) $\hat{\rho}$
200 +0.000+0.000 $0.070$ $0.964$ $0.968$ $1.040$
500 0.002-0.002 $0.045$ $0.946$ $0.946$ $1.014$
1000 0.001-0.001 $0.031$ $0.960$ $0.960$ $1.007$
2000 0.000-0.000 $0.023$ $0.958$ $0.958$ $1.003$
All coverage probabilities lie within $\pm 2$ Monte Carlo standard errors of the nominal 0.95 level (MC-SE $\approx 0.010$); fluctuations reflect simulation noise rather than systematic bias. The monotone decrease in $\hat{\rho}$ from 1.040 to 1.003 is the primary evidence of the strong-decay regime: as $n$ grows, $n\cdot\mathrm{Var}(R_{\rm rem})\to 0$ under $O_{P}(n^{-1/2})$ parametric nuisance convergence, the jackknife and sandwich converge to the same target, and no correction beyond the sandwich is needed. The jackknife CI is numerically identical to the sandwich CI once $\hat{\rho}\approx 1$, consistent with the sandwich characterization of Theorem 4.1. $V=1$ (no cross-fitting) is used to ensure the full-data and leave-one-out nuisance fits are structurally comparable; cross-fitting would be standard practice in applications but is not required for the strong-decay conclusion.
Under correctly specified parametric nuisance, $\hat{\rho}$ decreases monotonically from $1.040$ at $n=200$ to $1.003$ at $n=2000$, approaching $1.0$ as predicted by Theorem 4.1: when nuisance converges at the $O_{P}(n^{-1/2})$ parametric rate, $n\cdot\mathrm{Var}(R_{\rm rem})\to 0$ and the sandwich consistently estimates the total variance. The jackknife correction is unnecessary in this regime — both estimators target the same quantity and yield similar CIs once $\hat{\rho}\approx 1$. This contrasts with Table 1, where $\hat{\rho}$ remains above $1.14$ across all $J$ values under the SA-TMLE’s structural near-boundary remainder. Taken together, the two simulations illustrate the theoretical picture underlying Theorem 4.1: in this design, sandwich adequacy tracks the near-boundary condition on the remainder.
8 Discussion
We have developed a general theory of inference for asymptotically linear estimators when second-order remainder variance is non-negligible relative to the first-order influence-function variance. The central findings are:
  1. 1.
    In the near-boundary regime, the sandwich targets the influence-function variance $\sigma^{2}_{\mathrm{EIF}}/n$ rather than the total sampling variance, a gap identified structurally by the variance decomposition and characterized by Theorem 4.1.
  2. 2.
    The jackknife, pairs bootstrap, and HC correction each asymptotically target the total variance under the conditions established here, and yield valid confidence intervals. They differ in finite-sample accuracy, assumption requirements, and computational cost.
  3. 3.
    The variance ratio $\hat{\rho}$ provides an observable indicator of near-boundary operation that does not require knowledge of the remainder structure, as illustrated by the simulations.
  4. 4.
    In clustered data with shared random effects, the ICC amplifies the sandwich gap, a leading-order effect captured by Proposition 6.1 and consistent with the simulation evidence for SW-CRTs.
Connections and open questions.
Cross-fitting eliminates first-order empirical process terms but does not eliminate $\mathrm{Var}(R_{\mathrm{rem}})$, since the remainder is a $P_{0}$-expectation of a second-order product (Remark 3.2). The near-boundary gap could also be addressed analytically via a Bartlett correction, requiring an estimate of $c_{R}/\sigma^{2}_{\mathrm{EIF}}$; this is complementary to the resampling approach here. Open problems include non-asymptotic coverage bounds for the jackknife pivot and an analytic estimator for $c_{R}/\sigma^{2}_{\mathrm{EIF}}$ that avoids full jackknife refitting.
Conclusion.
The near-boundary regime identifies a structural gap between asymptotic linearity and valid first-order inference that is not resolved by standard product-rate conditions. The variance decomposition, sandwich characterization, and refined estimators here give a sharp account of this gap for general ALEs under the stated conditions. The most delicate assumptions concern the uniform LOO linearization and bootstrap moment consistency in highly adaptive nuisance regimes; extensions to these settings remain open.
Appendix A Proof of Proposition 6.1 (ICC Amplification)
We sketch the argument for the SA-TMLE DGP. Write the remainder as:
$R_{2}=\frac{1}{J}\sum_{j=1}^{J}\tilde{R}_{j},\quad\tilde{R}_{j}=\sum_{i,t}\bigl(\hat{\bar{Q}}_{Y}(S_{ijt},a,W_{ijt},t)-\bar{Q}_{Y}^{0}(\cdot)\bigr)\Bigl(\tfrac{\hat{g}_{\Delta}(\cdot)}{{g^{0}_{\Delta}}(\cdot)}-1\Bigr).$
The variance of $R_{2}$ is:
$\mathrm{Var}(R_{2})=\frac{1}{J^{2}}\sum_{j,j^{\prime}}\mathrm{Cov}(\tilde{R}_{j},\tilde{R}_{j^{\prime}})=\frac{1}{J}\mathrm{Var}(\tilde{R}_{j}),$
using cluster independence for $j\neq j^{\prime}$. Within cluster $j$, the observations share $b_{j}$, inducing correlation between the nuisance residuals $(\hat{\bar{Q}}_{Y}-\bar{Q}_{Y}^{0})(S_{ijt},\cdot)$ and $(\hat{\bar{Q}}_{Y}-\bar{Q}_{Y}^{0})(S_{ij^{\prime}t^{\prime}},\cdot)$ for $(i,t)\neq(i^{\prime},t^{\prime})$ in the same cluster. This between-observation correlation within a cluster contributes a term proportional to $\rho_{\mathrm{ICC}}\times m\times(\mathrm{average\;pairwise\;covariance\;of\;nuisance\;residuals})$, establishing (23).